### v1

# Branch structure and nonextensive thermodynamics of Kalb-Ramond-ModMax black holes: observational signatures

Abstract
We investigate a static, spherically symmetric black hole arising in Einstein gravity coupled to a Kalb-Ramond field and ModMax nonlinear electrodynamics, both of which are independently well motivated extensions of standard electrovacuum gravity. The solution depends, beyond mass and charge, on a Lorentz-violating parameter, a ModMax deformation parameter, and a discrete branch selector $\zeta=\pm 1$. We show that the ordinary branch admits extremal and non-extremal configurations, while the phantom branch generically supports a single-horizon geometry. Black-hole thermodynamics is analyzed within the Tsallis non-extensive framework, revealing branch-dependent stability and Joule-Thomson behavior. Weak gravitational lensing, photon propagation in plasma, and tidal forces are then studied, revealing clear optical and strong-field signatures that distinguish the two branches. In particular, the ordinary branch exhibits finite-radius tidal inversion, absent in the phantom sector. Our results demonstrate that the combined Kalb-Ramond and ModMax effects lead to a rich and observationally distinguishable black-hole phenomenology.
I Introduction
Black holes (BHs) occupy a central position in modern theoretical physics, providing a unique arena where gravity, quantum field theory, and high-energy physics intersect Giddings (2021); Barroso Varela and Rauch (2024). Within General Relativity (GR), the Schwarzschild and Reissner-Nordström (RN) solutions constitute the canonical descriptions of neutral and electrically charged BHs, respectively. Nevertheless, both theoretical considerations related to quantum gravity and a growing body of observational evidence from cosmology and astrophysics motivate the exploration of gravitational frameworks that extend or modify GR Saridakis and others (2021); Clifton et al. (2012). Such extensions often introduce additional fields, nonlinearities, or symmetry-breaking mechanisms, leading to nontrivial modifications of the causal structure, thermodynamics, and optical properties of BH spacetimes Nojiri and Odintsov (2011); Bambi (2017); Nashed and Saridakis (2019, 2020, 2022, 2023).
Among the most theoretically well-motivated extensions are those inspired by string theory and its low-energy effective descriptions. In this context, the Kalb-Ramond (KR) field, a rank-two antisymmetric tensor arising naturally in the bosonic and heterotic string spectra, provides a concrete realization of new gravitational degrees of freedom Lessa et al. (2020); Duan et al. (2024); Sucu and Sakalli (2025a). The associated three-form field strength can be interpreted either as spacetime torsion or as a vacuum condensate inducing Lorentz symmetry breaking (LSB) in the gravitational sector SenGupta and Sur (2001); Belinchón (2016). When coupled to Einstein gravity, the KR background modifies the spacetime geometry through an effective parameter $\ell$ that rescales the asymptotic structure while preserving spherical symmetry. The resulting BH geometries exhibit horizon and thermodynamic properties that differ from their GR counterparts, rendering them suitable probes of string-inspired effects in the strong-field regime Kao et al. (1996); Majumdar and SenGupta (1999).
In parallel, nonlinear electrodynamics (NED) has emerged as a powerful framework for addressing strong-field phenomena in gauge theories coupled to gravity Born and Infeld (1934); Ayon-Beato and Garcia (1998). Beyond the classic Born-Infeld construction Born and Infeld (1934), recent attention has focused on ModMax electrodynamics, which constitutes the unique one-parameter deformation of Maxwell theory preserving both conformal invariance and electric-magnetic duality Bandos et al. (2020, 2021). The ModMax Lagrangian introduces a dimensionless parameter $\gamma$ that interpolates continuously between the Maxwell limit and a genuinely nonlinear regime. When minimally coupled to gravity, ModMax electrodynamics gives rise to BH solutions with modified charge contributions and novel horizon structures Bandos et al. (2020); Sekhmani et al. (2025a). A distinctive feature of this theory is the emergence of two mathematically consistent branches: an ordinary branch ($\zeta=+1$), continuously connected to the RN geometry, and a phantom branch ($\zeta=-1$), in which the electromagnetic energy density effectively reverses sign, leading to qualitatively different physical behavior Eslam Panah et al. (2024).
The combined presence of LSB effects and nonlinear electromagnetic dynamics provides a particularly rich setting for BH physics Kostelecký and Li (2019); Sucu and Sakallı (2025b). In the Einstein-KR-ModMax framework, the spacetime geometry is controlled jointly by three deformation parameters: the KR parameter $\ell$, the ModMax nonlinearity parameter $\gamma$, and the branch selector $\zeta$. This interplay generates BH solutions with rescaled asymptotics, branch-dependent horizon structures, and modified thermodynamic and optical properties that distinguish them from both Schwarzschild and RN geometries Yang et al. (2023). Understanding how these different modifications combine and manifest in physical observables is essential for assessing their phenomenological relevance.
Black hole thermodynamics offers a particularly sensitive probe of the underlying gravitational dynamics. Since the formulation of the laws of BH mechanics and the discovery of Hawking radiation Bardeen et al. (1973); Hawking (1975), thermodynamic considerations have provided deep insights into the interplay between gravity, quantum theory, and statistical mechanics. However, the standard Boltzmann-Gibbs framework may be inadequate for gravitational systems characterized by long-range interactions and strong correlations Tsallis and Cirto (2013); Tsallis (1988). Non-extensive generalizations, most notably Tsallis entropy, introduce a deformation parameter $\delta$ that quantifies deviations from the area law and can significantly modify the thermodynamic behavior of BHs Tsallis and Cirto (2013); Saridakis et al. (2018); Basilakos et al. (2024). Within this framework, quantities such as internal energy, free energy, pressure, and heat capacity acquire nontrivial corrections, potentially altering phase structure and stability Abreu and Neto (2020); Nojiri et al. (2023); Sucu et al. (2026). The Joule-Thomson (JT) expansion further enriches this picture by providing a diagnostic of heating-cooling transitions in extended thermodynamic space Kruglov (2023); Liang et al. (2021); Sakallı et al. (2026); Zhang et al. (2024).
On the observational side, gravitational lensing remains one of the most direct probes of spacetime geometry in the strong-field regime. Since its first experimental confirmation Dyson et al. (1920), lensing has evolved into a precision tool for testing GR and constraining alternative theories Will (2014). The Gauss-Bonnet (GB) approach Gibbons and Werner (2008); Werner (2012) offers a coordinate-independent method for computing weak-field deflection angles and has been widely applied in BH spacetimes Ren et al. (2021); Sucu and Övgün (2024); Sucu (2026); Sucu and Sakalli (2025b); Wang et al. (2024). Moreover, realistic astrophysical environments are typically permeated by plasma, which modifies photon propagation through dispersive effects. In such settings, the photon sphere and shadow properties depend not only on the spacetime geometry but also on the plasma distribution, providing additional discriminants among competing gravitational models Perlick et al. (2015); Rogers (2015); Sucu et al. (2025).
Tidal forces constitute a complementary probe of the strong-field regime. Through the geodesic deviation equation, tidal effects encode local curvature information and govern the deformation of extended bodies approaching a BH Hobson et al. (2006); Varga and Grafarend (2018). In modified gravity scenarios, tidal force profiles may deviate substantially from the Schwarzschild case, potentially exhibiting enhanced magnitudes or sign reversals that influence tidal disruption processes Frolov and Novikov (2012); Burdge and others (2019); Poisson (2015).
Motivated by these considerations, in this work we perform a systematic investigation of Kalb-Ramond-ModMax BHs across both their ordinary and phantom branches. We analyze the spacetime structure and thermodynamics within the Tsallis entropy framework, including the JT expansion, and explore optical signatures through weak gravitational lensing and photon propagation in plasma environments. Finally, we examine tidal forces and identify branch-dependent features, including the conditions under which tidal inversion may occur. By combining geometric, thermodynamic, and observational diagnostics within a unified setting, we aim to clarify the physical implications of branch structure and non-extensive effects, and to identify signatures that could distinguish these BHs from standard GR solutions and other modified gravity models.
The paper is organized as follows. In Section II we present the KR-ModMax black hole solution, analyze its branch-dependent horizon structure, and derive the Hawking temperature. Section III develops the non-extensive thermodynamic description based on Tsallis entropy, while the Joule-Thomson expansion is examined in Section IV. In Section V we discuss the weak gravitational lensing and vacuum photon-sphere properties. Plasma effects on photon propagation and their impact on photon spheres are analyzed separately in Section VI. In Section VII we investigate tidal forces and branch-dependent deformations. Finally, Section VIII summarizes our results and outlines future directions.
II KR-ModMax BH Solution and Horizon Structure
We begin by introducing the gravitational setup and exact black hole solution arising from the combined presence of a Kalb-Ramond background and ModMax nonlinear electrodynamics. Particular emphasis is placed on the emergence of a branch structure, which plays a central role in shaping the causal, thermodynamic, and observational properties analyzed in subsequent sections.
II.1 Field equations, exact solution, and horizon structure
We consider static and spherically symmetric black hole geometries arising from Einstein gravity coupled simultaneously to a Kalb-Ramond (KR) three-form condensate and ModMax nonlinear electrodynamics. This framework combines Lorentz symmetry breaking (LSB) effects of string-theoretic origin with the unique conformally invariant deformation of Maxwell theory Sekhmani et al. (2025b); Bandos et al. (2020); Kruglov (2022). Owing to spherical symmetry, the spacetime metric can be written in the standard form
$ds^{2}=-f(r)\,dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}\,d\Omega^{2},$ (1)
where all geometric information is encoded in the lapse function $f(r)$, which captures both the KR-induced vacuum deformation and the nonlinear electromagnetic corrections Liu et al. (2024); Duan et al. (2024). The Einstein equations, supplemented by the modified Maxwell sector, reduce to a single ordinary differential equation for $f(r)$.
The KR field originates from an antisymmetric rank-two tensor $B_{\mu\nu}$ appearing in the massless sector of bosonic string theory. A nonvanishing condensate of its associated three-form field strength induces spontaneous breaking of local Lorentz invariance and rescales the effective gravitational and electromagnetic couplings through a dimensionless parameter $\ell$ Illuminati et al. (2021); Mukhopadhyaya et al. (2002). Moreover, ModMax electrodynamics constitutes the unique one-parameter nonlinear extension of Maxwell theory that preserves both conformal invariance and electric-magnetic duality Babaei-Aghbolagh et al. (2022); Bandos et al. (2020). The deviation from linear electrodynamics is governed by the parameter $\gamma$, with the Maxwell limit recovered as $\gamma\to 0$.
(a) Ordinary ModMax branch ($\zeta=+1$) for fixed $\ell=0.5$ and $\gamma=2$. Increasing electric charge $Q$ drives the geometry from non-extremal black holes (Q=0.3, 0.5Q=0.3,\,0.5) to the extremal configuration ($Q\simeq 0.93$), and finally to naked singularities (Q=1.0, 1.3Q=1.0,\,1.3).
(b) Phantom ModMax branch ($\zeta=-1$) for $\ell=0.5$ and $\gamma=2$. All configurations exhibit a single event horizon, with no extremal or naked solutions. The Schwarzschild limit ($Q=0$) is smoothly connected to the charged phantom geometries.
(c) Effect of the Kalb-Ramond parameter $\ell$ on the ordinary branch ($\zeta=+1$, $Q=0.5$, $\gamma=2$). Increasing $\ell$ shifts the event horizon inward and enhances the asymptotic normalization $f_{\infty}=(1-\ell)^{-1}$.
(d) Effect of the ModMax parameter $\gamma$ on the ordinary branch ($\zeta=+1$, $\ell=0.5$, $Q=0.7$). Larger $\gamma$ exponentially suppresses the effective electric charge through the factor $e^{-\gamma}$, continuously driving the geometry toward the Schwarzschild-KR limit.
Figure 1: Metric function $f(r)$ for KR-ModMax black holes in the ordinary ($\zeta=+1$) and phantom ($\zeta=-1$) branches. The horizontal dashed line $f(r)=0$ marks the locations of event horizons. Panel (a) illustrates the charge-driven transition from non-extremal to extremal and naked configurations in the ordinary branch. Panel (b) demonstrates the absence of extremal or naked solutions in the phantom sector. Panels (c) and (d) display the sensitivity of the geometry to the Kalb-Ramond parameter $\ell$ and the ModMax parameter $\gamma$, respectively.
In this setting, the modified Maxwell equations admit a purely radial electric field configuration $\Phi=\Phi(r)$, constrained by charge conservation. Direct integration yields the electrostatic potential
$\Phi(r)=\frac{Q\,e^{-\gamma}}{(1-\ell)\,r},$ (2)
where the factor $(1-\ell)^{-1}$ reflects the KR-induced rescaling of the effective electric coupling. Substituting Eq. (2) into the gravitational field equations and imposing asymptotic flatness leads to the exact metric function Sekhmani et al. (2025b)
$f(r)=\frac{1}{1-\ell}-\frac{2M}{r}+\frac{\zeta\,Q^{2}\,e^{-\gamma}}{(1-\ell)^{2}\,r^{2}},$ (3)
where $M$ is the ADM mass, $Q$ the electric charge, and $\zeta=\pm 1$ distinguishes between the ordinary ($\zeta=+1$) and phantom ($\zeta=-1$) branches of ModMax electrodynamics. While the ordinary branch corresponds to a positive electromagnetic energy density and continuously connects to the Reissner-Nordström solution, the phantom branch reverses the sign of the electromagnetic stress-energy contribution, leading to qualitatively different gravitational behavior.
The asymptotic structure of the spacetime follows from
$\lim_{r\to\infty}f(r)=\frac{1}{1-\ell},$ (4)
which shows that the KR parameter $\ell$ modifies the normalization of the timelike Killing vector at infinity. For $\ell<1$, the geometry remains asymptotically flat up to this constant rescaling, preserving the standard $1/r$ falloff of curvature invariants. Equivalently, the effective gravitational coupling is renormalized according to $G_{\text{eff}}=G/(1-\ell)$, with direct implications for orbital dynamics and gravitational lensing observables Masood (2024); Mukhopadhyaya et al. (2002).
In Fig. 1 we display the radial profile of the metric function $f(r)$ for four representative configurations. In particular, panel (a) illustrates the ordinary branch at fixed $(\ell,\gamma)=(0.5,2.0)$: as $Q$ increases from $0.3$ to $1.3$, the solution transitions from a non-extremal BH with two distinct horizons, through the extremal configuration at $Q\approx 0.93$, to a naked singularity for $Q\geq 1.0$. Panel (b) presents the phantom branch under identical KR-ModMax parameters, where all curves intersect the $f=0$ axis exactly once, confirming the single-horizon topology. Finally, panels (c) and (d) examine the sensitivity to $\ell$ and $\gamma$ respectively, demonstrating how these parameters shift horizon locations and modify the near-horizon geometry.
II.2 Branch dichotomy and physical interpretation
A distinctive feature of the Kalb-Ramond-ModMax black hole solution is the existence of two mathematically consistent branches, labeled by $\zeta=\pm 1$, which lead to qualitatively different spacetime structures and physical behavior. This branch dichotomy originates from the nonlinear electromagnetic sector and persists even in the presence of the Kalb-Ramond deformation, playing a central role in shaping the properties of the solution.
For the ordinary branch ($\zeta=+1$), the electromagnetic contribution partially counteracts the gravitational attraction, closely resembling the behavior familiar from the Reissner-Nordström geometry. As the electric charge approaches a critical value, the event and Cauchy horizons merge and the black hole reaches an extremal configuration with vanishing Hawking temperature. This limit is associated with the emergence of a degenerate horizon and signals a qualitative change in the near-horizon geometry. For larger charge values, the solution admits no horizons, giving rise to naked singularities.
On the other hand, the phantom branch ($\zeta=-1$) exhibits a markedly different structure. In this case, the effective electromagnetic contribution reinforces the gravitational field, eliminating the possibility of extremal configurations altogether. The spacetime admits a single event horizon for all values of the charge, and the Hawking temperature remains strictly positive throughout the parameter space. This absence of inner horizons implies a simpler causal structure and avoids the Cauchy horizon instabilities that typically afflict charged black holes in GR-like settings.
These geometric differences have direct physical implications. The reinforced gravitational potential characterizing the phantom branch modifies the location of the photon sphere and affects null geodesic propagation, while the ordinary branch retains features continuously connected to the Reissner-Nordström case. Similarly, the distinct horizon structures lead to different thermodynamic behavior, particularly near extremality, and influence tidal effects experienced by infalling matter. A geometric visualization of the equatorial spatial geometry, illustrating the single-horizon topology of the phantom branch, is presented in Appendix A.
The coexistence of ordinary and phantom branches within the same theoretical framework naturally raises the question of physical viability and observational discrimination. While the ordinary branch admits a smooth GR limit as $\ell\to 0$ and $\gamma\to 0$, the phantom branch represents a genuinely novel class of solutions with no direct counterpart in standard Einstein-Maxwell theory. In the following sections, we explore how this branch structure manifests in thermodynamic properties, optical signatures, and tidal forces, and assess whether these effects can provide meaningful criteria for distinguishing between the two branches.
(a) Ordinary ModMax branch ($\zeta=+1$). The electromagnetic contribution suppresses the Hawking temperature, producing Reissner-Nordström-like behavior. For small $\gamma$ (dark curves), $T_{H}$ approaches zero at finite $r_{h}$, signaling the onset of extremality. Increasing $\gamma$ exponentially weakens the effective charge through $e^{-\gamma}$, shifting the temperature curves upward and reducing the extremal horizon radius.
(b) Phantom ModMax branch ($\zeta=-1$). The sign-reversed electromagnetic sector enhances the Hawking temperature uniformly. All configurations remain strictly nonzero, precluding extremal limits. The temperature reaches $T_{H}\simeq 0.32$ at small $r_{h}$, approximately twice the maximum attained in the ordinary branch, reflecting reinforced thermal emission.
Figure 2: Hawking temperature $T_{H}$ as a function of the horizon radius $r_{h}$ for KR-ModMax black holes with fixed parameters $M=1.0$, $Q=0.5$, and $\ell=0.5$. The color scale denotes the ModMax parameter $\gamma\in[0,3]$. The ordinary branch exhibits a $\gamma$-dependent approach to extremality, whereas the phantom branch maintains strictly positive temperature throughout the entire parameter space.
II.3 Hawking temperature
The stationarity of the Kalb-Ramond-ModMax black hole spacetime ensures the existence of a timelike Killing vector associated with time translations,
$\xi^{\mu}=\left(\frac{\partial}{\partial t}\right)^{\mu},$ (5)
which gives rise to a conserved energy along timelike and null geodesics Hawking (1975). The surface gravity $\kappa$, characterizing the strength of the gravitational field at the event horizon, is defined through the Killing identity
$\nabla_{\nu}\left(\xi^{\mu}\xi_{\mu}\right)=-2\,\kappa\,\xi_{\nu},$ (6)
and remains constant over the horizon in accordance with the zeroth law of black hole mechanics Bardeen et al. (1973); Wald (1993).
For the static metric function (3), the surface gravity is obtained directly from the radial derivative of the lapse function evaluated at the outer horizon radius $r_{+}$,
$\kappa=\frac{1}{2}\,f^{\prime}(r)\Big|_{r=r_{+}}=\frac{M}{r_{+}^{2}}-\frac{\zeta\,Q^{2}\,e^{-\gamma}}{(1-\ell)^{2}\,r_{+}^{3}}.$ (7)
The relative sign of the two terms is controlled by the branch parameter $\zeta$, leading to qualitatively different thermal behavior in the ordinary and phantom branches.
Quantum particle production near the event horizon results in Hawking radiation, with temperature measured by asymptotic observers given by $T_{H}=\kappa/(2\pi)$ Hawking (1975). For the present geometry, this yields
$T_{H}=\frac{M}{2\pi r_{+}^{2}}-\frac{\zeta\,Q^{2}\,e^{-\gamma}}{2\pi(1-\ell)^{2}\,r_{+}^{3}}.$ (8)
This expression shows the combined influence of the Kalb-Ramond background, which rescales the effective couplings through $(1-\ell)^{-2}$, and the ModMax nonlinear electrodynamics, whose contribution is suppressed by the factor $e^{-\gamma}$ and whose sign is determined by $\zeta$.
In the ordinary branch ($\zeta=+1$), the electromagnetic contribution counteracts the gravitational term, closely mirroring the behavior of the Reissner-Nordström black hole. As the charge approaches the critical value $Q_{\text{ext}}=M\sqrt{(1-\ell)^{3}\,e^{\gamma}},$, the Hawking temperature vanishes and the horizons coincide, signaling the onset of an extremal configuration with degenerate surface gravity. This limit is associated with a qualitative change in the near-horizon geometry and marks the boundary of the semiclassical description Gibbons and Kallosh (1995); Hawking et al. (1994).
The phantom branch ($\zeta=-1$) exhibits a markedly different thermal structure. In this case, the electromagnetic term reinforces the gravitational contribution, leading to systematically higher temperatures and precluding the existence of extremal configurations. The Hawking temperature remains strictly positive throughout the parameter space, in agreement with the single-horizon causal structure discussed in the previous subsection.
Table 1: Horizon structure of KR-ModMax black holes with $M=1$. The ordinary branch ($\zeta=+1$) exhibits non-extremal, near-extremal, and naked singularity configurations depending on $Q$ relative to $Q_{\text{ext}}=M\sqrt{(1-\ell)^{3}e^{\gamma}}$. The phantom branch ($\zeta=-1$) admits only single-horizon black holes with $r_{+}$ increasing monotonically with charge. Configurations marked with correspond to the 3D embedding diagrams of Fig. 12 of Appendix A. Here $r_{\pm}$ denote outer/inner horizon radii and $T_{H}$ is the Hawking temperature.
$\zeta$ $\ell$ $\gamma$ $Q$ $_{+} rr_{-}$ $T_{H}$ Classification
Ordinary Branch ($\zeta=+1$) – Fig. 1(a)
+1+1 $0.5$ $2.0$ $0.3$ $0.975$ $0.025$ $0.159$ Non-extremal BH
+1+1 $0.5$ $2.0$ $0.5$ $0.927$ $0.073$ $0.158$ Non-extremal BH
+1+1 $0.5$ $2.0$ $0.7$ $0.843$ $0.157$ $0.154$ Non-extremal BH
+1+1 $0.5$ $2.0$ $0.9$ $0.675$ $0.325$ $0.122$ Near-extremal BH
+1+1 $0.5$ $2.0$ $1.0$ Naked Singularity
+1+1 $0.5$ $2.0$ $1.3$ Naked Singularity
Effect of KR Parameter $\ell$ ($\zeta=+1$, $Q=0.5$, $\gamma=2$) – Fig. 1(c)
+1+1 $0.0$ $2.0$ $0.5$ $1.983$ $0.017$ $0.040$ Non-extremal BH
+1+1 $0.2$ $2.0$ $0.5$ $1.573$ $0.027$ $0.062$ Non-extremal BH
+1+1 $0.4$ $2.0$ $0.5$ $1.151$ $0.049$ $0.110$ Non-extremal BH
+1+1 $0.5$ $2.0$ $0.5$ $0.927$ $0.073$ $0.158$ Non-extremal BH
Effect of ModMax Parameter $\gamma$ ($\zeta=+1$, $\ell=0.5$, $Q=0.7$) – Fig. 1(d)
+1+1 $0.5$ $0.0$ $0.7$ Naked Singularity
+1+1 $0.5$ $1.0$ $0.7$ Naked Singularity
+1+1 $0.5$ $2.0$ $0.7$ $0.843$ $0.157$ $0.154$ Non-extremal BH
+1+1 $0.5$ $3.0$ $0.7$ $0.912$ $0.088$ $0.157$ Non-extremal BH
Phantom Branch ($\zeta=-1$) – Fig. 1(b)
1-1 $0.5$ $2.0$ $0.0$ $1.000$ $0.159$ Single-horizon BH
1-1 $0.5$ $2.0$ $0.3$ $1.024$ $0.159$ Single-horizon BH
1-1 $0.5$ $2.0$ $0.5$ $1.064$ $0.159$ Single-horizon BH
1-1 $0.5$ $2.0$ $0.7$ $1.119$ $0.157$ Single-horizon BH
1-1 $0.5$ $2.0$ $0.9$ $1.185$ $0.155$ Single-horizon BH
3D Embedding Configurations ($\zeta=-1$) – Fig. 12
$1^{\dagger}$ $0.2$ $0.5$ $0.2$ $1.62$ $0.062$ Single-horizon BH
$1^{\dagger}$ $0.5$ $3.0$ $0.5$ $1.02$ $0.159$ Single-horizon BH
$1^{\dagger}$ $0.7$ $1.0$ $0.8$ $1.24$ $0.104$ Single-horizon BH
$1^{\dagger}$ $0.4$ $2.0$ $0.1$ $1.20$ $0.110$ Single-horizon BH
The branch-dependent behavior of the Hawking temperature is illustrated in Fig. 2, where $T_{H}$ is shown as a function of the horizon radius for representative values of the ModMax parameter $\gamma$. In the ordinary branch [Fig. 2(a)], increasing $\gamma$ suppresses the effective charge contribution and progressively restores Schwarzschild-like behavior. By contrast, in the phantom branch [Fig. 2(b)], the temperature curves are shifted upward and exhibit a smoother dependence on $\gamma$, reflecting the absence of extremality.
Table 1 summarizes the resulting horizon structure across representative parameter combinations for both branches, explicitly illustrating the transition between non-extremal, extremal, and horizonless configurations. The numerical values reported in Table 1 are fully consistent with the aforementioned trends and further reveal the enhanced thermal robustness of the phantom branch against charge variations.
III Tsallis Entropy and KR-ModMax BH Thermodynamics
Black hole thermodynamics provides a sensitive probe of the microscopic structure of gravitational theories. In the presence of nonlinear matter sectors and Lorentz symmetry breaking backgrounds, departures from standard extensive thermodynamics may naturally arise. In this section, we adopt a non-extensive framework based on Tsallis entropy in order to examine how such effects modify the thermodynamic behavior of Kalb-Ramond-ModMax black holes and how the resulting properties depend on the underlying branch structure.
III.1 Tsallis entropy, internal energy, and free energy
Within the Tsallis framework, the entropy of a black hole with event-horizon radius $r_{h}$ is taken to be Tsallis and Cirto (2013)
$S_{T}=\left(\pi r_{h}^{2}\right)^{\delta},$ (9)
where $\delta$ denotes the non-extensivity parameter. The standard Bekenstein-Hawking entropy $S_{\text{BkH}}=\pi r_{h}^{2}$ (in units $G=\hbar=c=k_{B}=1$) is recovered for $\delta=1$ Bekenstein (1980); Hawking (1975). Values $\delta\neq 1$ encode deviations from extensivity, reflecting the possible presence of long-range correlations, horizon fluctuations, or an underlying fractal microstructure associated with quantum gravitational effects Biró and Czinner (2013). In what follows, we focus on $\delta\geq 1$, which ensures physical behavior and avoids pathological divergences.
The internal energy associated with the Tsallis entropy is obtained from the first law of black hole thermodynamics,
$E_{T}=\int T_{H}\,dS_{T},$ (10)
where the non-extensive nature of the entropy introduces a $\delta$-dependent weighting through $dS_{T}/dr_{h}=2\delta\,\pi^{\delta}\,r_{h}^{2\delta-1}$. Substituting the Hawking temperature given in Eq. (8) and performing the integration yields
$E_{T}=\frac{\mathcal{A}}{2r_{h}^{3}(1-\ell)^{2}e^{\gamma}\pi(2\delta-1)(4\delta-3)},$ (11)
with the numerator
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\mathcal{A}=\delta\,r_{h}^{4\delta}\pi^{2\delta}\Big[4Me^{\gamma}\delta\ell^{2}r_{h}-8Me^{\gamma}\delta\ell r_{h}$
$\ \ \ \ \ -3Me^{\gamma}\ell^{2}r_{h}+4Me^{\gamma}\delta r_{h}+6Me^{\gamma}\ell r_{h}$
$\ \ \ \ \ -3Me^{\gamma}r_{h}-4Q^{2}\delta\zeta+2\zeta Q^{2}\Big].$ (12)
The denominators $(2\delta-1)$ and $(4\delta-3)$ formally signal divergences at $\delta=1/2$ and $\delta=3/4$, respectively, however these values lie outside the physically relevant regime considered here.
The behavior of the internal energy as a function of the horizon radius and the non-extensivity parameter is illustrated in Fig. 3 for both ModMax branches. In the ordinary branch (Fig. 3(a)), $E_{T}$ increases monotonically with $r_{h}$ and $\delta$, reflecting the standard competition between gravitational and electromagnetic contributions. In the phantom branch (Fig. 3(b)), the internal energy exhibits a similar qualitative dependence but is systematically enhanced. This increase originates from the sign reversal $\zeta=-1$, which effectively converts the electromagnetic contribution into an additional heating source, leading to a larger accumulated thermal energy for fixed horizon size.
(a) Ordinary ModMax branch ($\zeta=+1$). The Tsallis internal energy $E_{T}$ increases monotonically with the horizon radius $r_{h}$, displaying smooth scaling across the non-extensivity range $\delta\in[1,2]$. Charge effects suppress the short-distance contribution, leading to a gradual approach toward Schwarzschild-KR behaviour at large radii.
(b) Phantom ModMax branch ($\zeta=-1$). The internal energy exhibits systematically steeper growth compared to the ordinary branch. The sign-reversed electromagnetic sector amplifies the near-horizon contribution, enhancing both the magnitude of $E_{T}$ and its sensitivity to the non-extensivity parameter $\delta$.
Figure 3: Tsallis internal energy $E_{T}$ as a function of the horizon radius $r_{h}$ for KR-ModMax black holes with fixed parameters $M=1.0$, $Q=0.5$, $\ell=0.5$, and $\gamma=2.0$. The color scale denotes the non-extensivity parameter $\delta\in[1,2]$. Across the entire parameter space, the phantom branch maintains higher internal energy, establishing a clear thermodynamic hierarchy between the two ModMax sectors.
(a) Ordinary ModMax branch ($\zeta=+1$). The Tsallis Helmholtz free energy $F_{T}$ decreases monotonically with increasing horizon radius $r_{h}$, approaching zero asymptotically. Larger non-extensivity parameter $\delta$ deepens the free-energy well at small $r_{h}$, indicating a higher thermodynamic cost for compact configurations and a tendency toward instability near the horizon.
(b) Phantom ModMax branch ($\zeta=-1$). The free energy attains systematically larger magnitudes at small $r_{h}$ and exhibits a steeper radial dependence. The enhanced separation between $\delta$-curves reflects a stronger sensitivity to non-extensive corrections, signaling a more pronounced thermodynamic instability in the phantom sector.
Figure 4: Tsallis Helmholtz free energy $F_{T}$ as a function of the horizon radius $r_{h}$ for KR-ModMax black holes with the same parameters as Fig. 3. While both branches show monotonic decay with increasing $r_{h}$, the phantom branch maintains a larger free-energy magnitude throughout, reinforcing the thermodynamic hierarchy inferred from the internal energy and anticipating the branch-dependent stability behavior discussed in subsequent sections.
A complementary thermodynamic quantity is the Helmholtz free energy, which characterizes the maximum extractable work at fixed temperature and provides insight into global stability. Within the Tsallis framework, it is obtained through the Legendre transformation
$F_{T}=-\int S_{T}\,dT_{H}.$ (13)
Performing the integration yields
$F_{T}=\frac{\mathcal{K}}{r_{h}^{3}(1-\ell)^{2}\pi(2\delta-1)(4\delta-3)},$ (14)
where
$\!\!\!\!\!\!\!\!\mathcal{K}=(1-\ell)^{2}\left(\delta-\frac{3}{4}\right)\pi^{2\delta}M\,r_{h}^{4\delta+1}$
$\ \ -\left(\delta-\frac{1}{2}\right)\left[(4\delta-3)\left(-4\zeta Q^{2}e^{-\gamma}+4Mr_{h}(1-\ell)^{2}\right)\right$.
$\ \ \left.+\frac{3e^{-\gamma}\pi^{2\delta}r_{h}^{4\delta}Q^{2}\zeta}{2}\right].$ (15)
The behavior of the Helmholtz free energy is displayed in Fig. 4. In the ordinary branch, $F_{T}$ decreases monotonically with the horizon radius, with increasing $\delta$ enhancing the preference for larger, thermodynamically stable configurations. In the phantom branch, the free energy exhibits a similar qualitative trend but with systematically larger magnitude and steeper descent, indicating that non-extensive effects couple more strongly when the electromagnetic contribution reinforces the gravitational field Al-Badawi et al. (2025).
III.2 Thermodynamic pressure and equation of state
In the extended formulation of black hole thermodynamics, macroscopic thermodynamic variables can be introduced through appropriate conjugate pairs, allowing one to explore phase structure beyond the standard entropy-temperature description Kubiznak and Mann (2012). Within the present Tsallis framework, we define an effective thermodynamic pressure through the Helmholtz free energy, while retaining the geometric interpretation of the thermodynamic volume as
$V=\frac{4\pi r_{h}^{3}}{3}.$ (16)
The pressure is then obtained as
$P_{T}=-\frac{\partial F_{T}}{\partial V},$ (17)
which encodes how non-extensive effects and matter-sector modifications reshape the effective equation of state. Hence, carrying out the differentiation yields
$P_{T}=\frac{\mathcal{D}}{16(1-\ell)^{2}r_{h}^{6}\pi^{2}},$ (18)
with
$\mathcal{D}=-2M\pi^{2\delta}(1-\ell)^{2}r_{h}^{4\delta+1}+3e^{-\gamma}\pi^{2\delta}r_{h}^{4\delta}Q^{2}\zeta\\ +6\zeta Q^{2}e^{-\gamma}-4Mr_{h}(1-\ell)^{2}.$ (19)
As we can see, the resulting pressure is negative throughout the parameter space, reflecting the attractive nature of the effective interaction governing the black hole thermodynamics in this setting.
The behavior of $P_{T}$ as a function of the horizon radius is illustrated in Fig. 5. In the ordinary branch (Fig. 5(a)), the pressure decreases smoothly in magnitude as $r_{h}$ increases, with larger values of the non-extensivity parameter $\delta$ systematically reducing the absolute pressure. In the phantom branch (Fig. 5(b)), a similar qualitative trend is observed, although the magnitude of the pressure decreases more rapidly at small horizon radii, reflecting the enhanced role of the electromagnetic contribution when $\zeta=-1$. At sufficiently large $r_{h}$, both branches converge as the mass term dominates over charge-dependent corrections.
This effective equation of state provides the thermodynamic background for the Joule-Thomson expansion discussed in the following section, where heating and cooling regimes are analyzed in terms of the response of the Hawking temperature to pressure variations.
(a) Ordinary ModMax branch ($\zeta=+1$). The Tsallis thermodynamic pressure $P_{T}$ remains negative throughout the explored parameter space and decreases smoothly with increasing horizon radius $r_{h}$. Larger non-extensivity parameter $\delta$ uniformly reduces the pressure magnitude, indicating that non-extensive entropy softens the effective equation of state and moderates the response to horizon shrinkage.
(b) Phantom ModMax branch ($\zeta=-1$). The pressure exhibits a similar negative range but displays a steeper radial dependence at small $r_{h}$. The sign-reversed electromagnetic sector enhances the sensitivity of $P_{T}$ to near-horizon physics, foreshadowing stronger branch dependence in pressure-driven processes such as Joule-Thomson expansion.
Figure 5: Tsallis thermodynamic pressure $P_{T}$ as a function of the horizon radius $r_{h}$ for KR-ModMax black holes. The negative pressure indicates that, in this asymptotically flat setup, the effective cosmological contribution acts as a tension rather than a true pressure. Although both branches converge at large $r_{h}$ where the mass term dominates, their distinct radial behavior at small $r_{h}$ plays a crucial role in shaping the Joule-Thomson coefficient and the cooling-heating structure discussed in Sec. IV.
III.3 Heat capacity and local stability
(a) Ordinary ModMax branch ($\zeta=+1$). The heat capacity $C_{V}$ remains negative throughout the explored range $r_{h}\in[1.0,\,1.25]$, indicating local thermodynamic instability. As $r_{h}$ increases, $C_{V}$ approaches zero from below, signaling proximity to a Davies-type transition where the dominant instability weakens. The stratification with respect to the non-extensivity parameter $\delta$ reflects the increasing weight of entropy corrections near the transition region.
(b) Phantom ModMax branch ($\zeta=-1$). The heat capacity exhibits the same qualitative instability pattern but with systematically reduced magnitude. The absence of a zero-crossing confirms that no stable phase is realized, while the smoother approach toward $C_{V}\to 0^{-}$ indicates a weaker near-critical response compared to the ordinary branch.
Figure 6: Heat capacity $C_{V}$ as a function of the horizon radius $r_{h}$ for KR-ModMax black holes within Tsallis non-extensive thermodynamics. Negative values throughout indicate local thermodynamic instability for both branches. The branch-dependent approach toward $C_{V}=0^{-}$ at larger $r_{h}$ plays a decisive role in shaping the Joule-Thomson cooling-heating behavior discussed in Sec. IV. The color scale corresponds to the non-extensivity parameter $\delta\in[1.0,\,2.0]$.
Local thermodynamic stability of the black hole is determined by the heat capacity at constant volume, defined as
$C_{V}=T_{H}\left(\frac{\partial S_{T}}{\partial T_{H}}\right)_{V},$ (20)
where a positive heat capacity corresponds to stability under thermal fluctuations, while a negative value signals local thermodynamic instability. Within the Tsallis framework adopted here, the heat capacity for the Kalb-Ramond-ModMax black hole can be evaluated analytically, yielding
$C_{V}=-\frac{2\pi^{2\delta}\left[Mr_{h}(1-\ell)^{2}e^{\gamma}-\zeta Q^{2}\right]r_{h}^{4\delta}\delta}{2Mr_{h}(1-\ell)^{2}e^{\gamma}-3\zeta Q^{2}}.$ (21)
The denominator of Eq. (21) vanishes at $r_{h}=3\zeta Q^{2}e^{-\gamma}/[2M(1-\ell)^{2}]$, indicating a divergence of $C_{V}$ that separates thermodynamically distinct regimes.
The sign and magnitude of the heat capacity are strongly branch dependent. In the ordinary branch ($\zeta=+1$), the electromagnetic contribution counteracts the gravitational term, leading to negative heat capacity over a broad range of horizon radii. In the phantom branch ($\zeta=-1$), the same contribution reinforces gravity, modifying the location and strength of the divergence but not altering the overall sign of $C_{V}$ in the near-horizon region.
The behavior of the heat capacity is illustrated in Fig. 6 for representative parameter choices. In both branches, $C_{V}$ remains negative over the displayed range of horizon radii, indicating local thermodynamic instability in the near-horizon regime. The ordinary branch (Fig. 6(a)) exhibits a monotonic increase of $C_{V}$ toward zero as $r_{h}$ grows, with the non-extensivity parameter $\delta$ producing a systematic stratification of the curves. The phantom branch (Fig. 6(b)) displays a similar qualitative trend, with slightly reduced magnitude, reflecting the reinforcing role of the electromagnetic sector.
The approach of $C_{V}$ toward zero at larger horizon radii suggests proximity to a Davies-type transition Shahzad and Nosheen (2022), beyond which locally stable configurations may arise. However, within the parameter range explored here, both branches remain thermodynamically unstable at the local level, a feature that complements the global stability analysis based on the Helmholtz free energy and motivates the extended thermodynamic investigation presented in the next section.
IV Joule-Thomson expansion
The Joule-Thomson (JT) expansion provides a valuable diagnostic of thermodynamic response by examining the evolution of the Hawking temperature under isenthalpic processes. Within the extended thermodynamic framework developed above, it offers insight into how non-extensive effects and branch structure influence heating and cooling behavior beyond equilibrium considerations.
IV.1 Joule-Thomson coefficient and inversion curve
The JT coefficient is defined as Aydiner et al. (2025)
$\mu_{J}=\left(\frac{\partial T_{H}}{\partial P_{T}}\right),$ (22)
with $\mu_{J}>0$ corresponding to cooling during expansion and $\mu_{J}<0$ indicating heating. The inversion curve, defined by the condition $\mu_{J}=0$, separates these regimes in the $(P_{T},T_{H})$ plane Wang et al. (2025).
For the Kalb-Ramond-ModMax black hole, explicit evaluation yields
$\mu_{J}=\frac{4\left[Mr_{h}(1-\ell)^{2}e^{\gamma}-\frac{3Q^{2}}{2}\right]\pi r_{h}^{3}}{\mathcal{B}},$ (23)
where
$\mathcal{B}=2M\pi^{2\delta}\left(\delta-\frac{5}{4}\right)(1-\ell)^{2}e^{\gamma}r_{h}^{4\delta+1}\\ -3\pi^{2\delta}\left(\delta-\frac{3}{2}\right)Q^{2}r_{h}^{4\delta}-5Mr_{h}(1-\ell)^{2}e^{\gamma}+9Q^{2}.$ (24)
The numerator of Eq. (23) vanishes at $r_{h}=3Q^{2}e^{-\gamma}/[2M(1-\ell)^{2}]$, defining the inversion radius at which the JT coefficient changes sign. The location of this inversion point depends explicitly on the ModMax parameter $\gamma$, the KR deformation $\ell$, and the non-extensivity parameter $\delta$, highlighting the sensitivity of the JT process to the underlying microscopic structure.
Figure 7: Joule-Thomson (JT) coefficient $\mu_{J}$ as a function of the horizon radius $r_{h}$ for the ordinary-branch KR-ModMax black hole ($M=1$, $Q=0.5$, $\zeta=+1$, $\gamma=2$, $\ell=0.5$). The color scale corresponds to the non-extensivity parameter $\delta\in[1.0,\,2.0]$. Divergences of $\mu_{J}$ occur at $r_{h}\approx 1.2$ and $r_{h}\approx 2.7$-$3.0$, signaling inversion points that separate heating ($\mu_{J}<0$) from cooling ($\mu_{J}>0$) regimes during isenthalpic expansion. Increasing $\delta$ shifts the inversion radii toward larger $r_{h}$, demonstrating that non-extensive entropy systematically delays the onset of the cooling phase.
IV.2 Cooling-heating transitions and branch dependence
The behavior of the JT coefficient as a function of the horizon radius is illustrated in Fig. 7 for the ordinary branch. The coefficient exhibits pronounced divergences associated with the vanishing of the denominator $\mathcal{B}$, separating distinct thermodynamic response regimes. In the near-horizon region, $\mu_{J}$ is negative, indicating heating during expansion as electromagnetic contributions dominate. At intermediate radii, $\mu_{J}$ becomes positive, signaling a cooling regime in which the mass term governs the thermodynamic response. For sufficiently large horizon radii, the JT coefficient approaches small positive values and gradually stabilizes.
The non-extensivity parameter $\delta$ plays a crucial role in shaping the inversion structure. Increasing $\delta$ systematically shifts the inversion points toward larger horizon radii, implying that non-extensive corrections delay the onset of the cooling regime by enhancing the entropy contribution in the inversion condition. This behavior reflects the fact that Tsallis entropy amplifies the weight of large-horizon configurations in the thermodynamic response.
Although the figure focuses on the ordinary branch, the analytic structure of Eq. (23) indicates that the branch parameter $\zeta$ influences both the location of divergences and the extent of cooling and heating regions. In particular, the reinforced electromagnetic contribution characterizing the phantom branch is expected to modify the balance between heating and cooling phases, in line with the enhanced thermal behavior observed in the previous sections. Overall, the JT expansion provides a sensitive probe of how non-extensive thermodynamics and branch structure jointly affect the response properties of Kalb-Ramond-ModMax black holes Kruglov (2022).
V Gravitational lensing and photon spheres
Gravitational lensing provides a direct and geometry-driven probe of spacetime structure around compact objects, allowing deviations from general relativity to be tested without detailed assumptions about the nature of the light source Cardoso and Pani (2019); Chakraborty and SenGupta (2017). Since the first experimental confirmation of light deflection by the Sun, lensing has evolved into a precision tool for constraining modified gravity models, including scenarios with Lorentz symmetry breaking or nonlinear electromagnetic sectors Partanen and Tulkki (2025). In the weak-field regime, where the photon trajectory remains far from the event horizon, small deviations from the Schwarzschild prediction can accumulate into potentially observable signatures relevant to astrophysical lensing systems Tsukamoto et al. (2014).
In the present context, weak gravitational lensing offers a sensitive diagnostic of the Kalb-Ramond-ModMax black hole. Although both the Kalb-Ramond condensate and the ModMax nonlinearity primarily affect the near-horizon geometry, their influence propagates to large distances through the optical structure of spacetime. To capture these effects in a coordinate-independent manner, we employ the Gauss-Bonnet (GB) theorem approach Gibbons and Werner (2008); Sucu and Sakalli (2025c); Sucu and Sakallı (2025c), which interprets the deflection of light as a global geometric effect governed by the integrated Gaussian curvature of an associated optical manifold.
V.1 Optical geometry and Gauss-Bonnet method
In static and spherically symmetric spacetimes, null geodesics can be equivalently described as spatial geodesics on an effective two-dimensional Riemannian manifold known as the optical geometry ABRAMOWICZ et al. (1988); Perlick (2006). Restricting the Kalb-Ramond-ModMax metric (1) to the equatorial plane ($\theta=\pi/2$), the corresponding optical line element takes the form
$d\sigma^{2}=\frac{dr^{2}}{f^{2}(r)}+\frac{r^{2}}{f(r)}\,d\phi^{2},$ (25)
where the lapse function $f(r)$ is given in Eq. (3). The conformal factor $1/f(r)$ encodes gravitational redshift effects and ensures that extremizing the optical path length reproduces the null geodesic equations via Fermat’s principle Wei et al. (2015).
The intrinsic curvature of the optical manifold is characterized by the Gaussian curvature $K$, which in two dimensions is fully determined by a single independent component of the Riemann tensor Sucu and Sakallı (2025a), namely
$K=\frac{R_{r\phi r\phi}}{\det(g_{\rm opt})}.$ (26)
In the weak-field region ($r\gg r_{h}$), we expand the metric function asymptotically and retain terms up to $\mathcal{O}(r^{-5})$. This yields the following expression for the Gaussian curvature:
$\!\!\!\!\!K\simeq\frac{3\zeta Q^{2}}{r^{4}(1-\ell)^{4}e^{\gamma}}-\frac{3\zeta Q^{2}\ell}{r^{4}(1-\ell)^{4}e^{\gamma}}+\frac{2M\ell^{3}}{r^{3}(1-\ell)^{4}}$
$\ \ \ \ -\frac{6M\ell^{2}}{r^{3}(1-\ell)^{4}}+\frac{6M\ell}{r^{3}(1-\ell)^{4}}-\frac{2M}{r^{3}(1-\ell)^{4}}$
$\ \ \ \ +\frac{3M^{2}\ell^{4}}{r^{4}(1-\ell)^{4}}-\frac{12M^{2}\ell^{3}}{r^{4}(1-\ell)^{4}}+\frac{18M^{2}\ell^{2}}{r^{4}(1-\ell)^{4}}$
$\ \ \ \ -\frac{12M^{2}\ell}{r^{4}(1-\ell)^{4}}+\frac{3M^{2}}{r^{4}(1-\ell)^{4}}.$ (27)
The leading $r^{-3}$ contributions originate from the mass sector, while the $r^{-4}$ terms encode both second-order mass corrections and the charge-dependent ModMax contributions, controlled by the branch parameter $\zeta$ and the suppression factor $e^{-\gamma}$. The Kalb-Ramond parameter $\ell$ enters universally through the prefactor $(1-\ell)^{-4}$, amplifying all curvature contributions as $\ell$ approaches unity.
The Gauss-Bonnet theorem provides a direct link between this intrinsic curvature and the total deflection angle experienced by a light ray. Applied to a non-singular domain $\mathcal{D}_{\infty}$ exterior to the photon trajectory, the theorem yields the coordinate-independent expression Gibbons and Werner (2008)
$\beta=-\iint_{\mathcal{D}_{\infty}}K\,dS,$ (28)
where $dS$ denotes the surface element of the optical manifold. In the weak-lensing approximation, the photon path is well approximated by the unperturbed straight-line trajectory $r=b/\sin\phi$, allowing the integral to be evaluated analytically. The resulting deflection angle, together with its branch-dependent structure and limiting behavior, is discussed in the following subsection.
V.2 Weak-field deflection angle and branch-dependent signatures
Within the Gauss-Bonnet framework, the total deflection angle experienced by a light ray with impact parameter $b$ is determined by the surface integral of the Gaussian curvature over the exterior optical domain $\mathcal{D}_{\infty}$, namely
$\beta=-\iint_{\mathcal{D}_{\infty}}K\,dS,$ (29)
where the integration region is bounded by the photon trajectory and an asymptotic reference curve Gibbons and Werner (2008). In the weak-field regime, the photon path is well approximated by the unperturbed straight-line trajectory $r=b/\sin\phi$, which allows the integral to be evaluated analytically.
Carrying out the integration using the asymptotic expansion of the Gaussian curvature given in Eq. (27), we obtain the vacuum deflection angle for the Kalb-Ramond-ModMax black hole,
$\!\!\!\!\!\!\beta_{\rm vac}\approx\frac{4M}{b(1-\ell)^{4}}-\frac{12M\ell}{b(1-\ell)^{4}}+\frac{12M\ell^{2}}{b(1-\ell)^{4}}-\frac{4M\ell^{3}}{b(1-\ell)^{4}}$
$\!\!\!\!\!\!+\frac{15\pi M^{2}}{4b^{2}(1-\ell)^{4}}-\frac{3\pi M^{2}\ell}{2b^{2}(1-\ell)^{4}}+\frac{3\pi M^{2}\ell^{3}}{2b^{2}(1-\ell)^{4}}-\frac{3\pi M^{2}\ell^{4}}{4b^{2}(1-\ell)^{4}}$
$\!\!\!\!\!\!-\frac{3\pi Q^{2}\zeta}{4b^{2}(1-\ell)^{4}e^{\gamma}}+\frac{3\pi Q^{2}\zeta\ell}{4b^{2}(1-\ell)^{4}e^{\gamma}}-\frac{4MQ^{2}\zeta}{3b^{3}(1-\ell)^{4}e^{\gamma}}$
$\!\!\!\!\!\!+\frac{8MQ^{2}\ell^{2}\zeta}{3b^{3}(1-\ell)^{4}e^{\gamma}}-\frac{4MQ^{2}\ell\zeta}{3b^{3}(1-\ell)^{4}e^{\gamma}}+\mathcal{O}(M^{3},1/b^{4}).$ (30)
This expression naturally decomposes into three classes of contributions: (i) mass terms proportional to $M/b$ and $M^{2}/b^{2}$, (ii) Kalb-Ramond corrections controlled by powers of $\ell$ and the universal amplification factor $(1-\ell)^{-4}$, and (iii) charge-dependent ModMax terms proportional to $Q^{2}\zeta$, whose magnitude is suppressed by $e^{-\gamma}$ and whose sign is fixed by the branch parameter $\zeta$.
Several physically relevant limits follow immediately. In the Schwarzschild limit $(\ell\to 0,\gamma\to 0,Q\to 0)$, Eq. (30) reduces to
$\beta_{\rm Sch}=\frac{4M}{b}+\frac{15\pi M^{2}}{4b^{2}}+\mathcal{O}(b^{-3}),$ (31)
recovering the standard general-relativistic result Epstein and Shapiro (1980); Weinberg (2013). For the ordinary branch $(\zeta=+1)$ with $\ell\to 0$ and $\gamma\to 0$, the expression reduces to the Reissner-Nordström deflection angle, i.e.
$\beta_{\rm RN}=\frac{4M}{b}+\frac{15\pi M^{2}}{4b^{2}}-\frac{3\pi Q^{2}}{4b^{2}}-\frac{4MQ^{2}}{3b^{3}}+\mathcal{O}(b^{-4}),$ (32)
where the electromagnetic contribution partially counteracts gravitational focusing.
On the other hand, in the phantom branch $(\zeta=-1)$ the sign reversal of the charge sector transforms the electromagnetic contribution into an additional focusing mechanism. Even in the limit $\ell\to 0$ and $\gamma\to 0$, the deflection angle becomes
$\beta_{\rm phantom}=\frac{4M}{b}+\frac{15\pi M^{2}}{4b^{2}}+\frac{3\pi Q^{2}}{4b^{2}}+\frac{4MQ^{2}}{3b^{3}}+\mathcal{O}(b^{-4}),$ (33)
which has no analogue in standard general relativity. In this sector, both mass and charge contribute constructively to light bending, leading to systematically larger deflection angles.
(a) Ordinary branch ($\zeta=+1$).
(b) Phantom branch ($\zeta=-1$).
Figure 8: Vacuum deflection angle $\beta_{\rm vac}$ as a function of the impact parameter $b$ for KR-ModMax black holes with fixed $M=1$, $Q=0.5$, and $\gamma=2$. The color scale indicates the Kalb-Ramond parameter $\ell\in[0,0.5]$. Panel (a) shows the ordinary branch ($\zeta=+1$), where the electromagnetic contribution partially counteracts gravitational focusing, leading to suppressed deflection at small $b$. Increasing $\ell$ enhances the mass-dominated terms through the $(1-\ell)^{-4}$ amplification. Panel (b) displays the phantom branch ($\zeta=-1$), where the sign reversal of the electromagnetic sector transforms the charge contribution into an additional focusing mechanism. Consequently, the deflection angle remains strictly positive and is notably enhanced at small $b$, exceeding the ordinary-branch values by up to an order of magnitude as $\ell$ increases. This qualitative distinction serves as an observational discriminator between the two electromagnetic sectors.
These qualitative differences are evident in the numerical results shown in Fig. 8, where the deflection angle is plotted for varying Kalb-Ramond parameter $\ell\in[0,0.5]$ at fixed $\gamma=2$. In the ordinary branch, the competition between mass-induced focusing and charge-induced defocusing produces smaller deflection angles, particularly at low $\ell$. As $\ell$ increases, the universal amplification factor $(1-\ell)^{-4}$ enhances all contributions, leading to progressively larger deflection at all impact parameters. In the phantom branch, the deflection angle remains strictly positive and is substantially enhanced relative to the ordinary branch, exceeding the latter by more than an order of magnitude at small $b$ for larger $\ell$ values. This pronounced enhancement provides a clear geometric manifestation of the branch dichotomy.
From an observational perspective, these branch-dependent signatures translate into potentially measurable effects. Phantom configurations would produce significantly larger Einstein ring radii and enhanced magnification in strong-lensing systems, while the $\ell$-dependence of the weak-field deflection could imprint small but systematic astrometric shifts in high-precision measurements of stellar orbits near supermassive black holes. Together with the thermodynamic and photon-sphere properties discussed in previous sections, gravitational lensing offers a complementary channel for discriminating between the ordinary and phantom Kalb-Ramond-ModMax black hole branches.
VI Plasma effects on photon propagation
In realistic astrophysical environments, black holes are rarely isolated systems. Instead, they are typically surrounded by plasma-rich media such as accretion disks, stellar winds, or diffuse interstellar matter, all of which modify photon propagation through dispersive effects Broderick and Loeb (2009). In contrast to vacuum null geodesics, light rays propagating in plasma experience frequency-dependent refraction, leading to shifts in photon trajectories, photon-sphere locations, and shadow properties.
For the Kalb-Ramond-ModMax black hole, incorporating plasma effects is therefore essential for bridging the gap between theoretical predictions and observational applications, particularly in the context of horizon-scale imaging and lensing measurements by instruments such as the Event Horizon Telescope Akiyama and others (2019). In this section, we analyze photon dynamics in a dispersive plasma medium and derive the modified conditions for circular photon orbits that define the photon sphere.
VI.1 Photon dynamics and photon spheres in plasma
Photon propagation in a plasma medium can be described using an effective Hamiltonian formalism that incorporates the dispersive response of the medium Crisnejo and Gallo (2018). For a static plasma distribution in the Kalb-Ramond-ModMax spacetime, the Hamiltonian governing photon motion takes the form
$H=\frac{1}{2}\left[g^{\alpha\beta}p_{\alpha}p_{\beta}+(\omega_{p}^{2}-1)(p_{\alpha}u^{\alpha})^{2}\right],$ (34)
where $p_{\alpha}$ denotes the photon four-momentum, $u^{\alpha}$ is the plasma four-velocity, and $\omega_{p}$ is the local plasma frequency. Assuming a static plasma configuration, the four-velocity is given by $u^{\alpha}=f^{-1/2}\delta^{\alpha}_{t}$, where $f(r)$ is the KR-ModMax lapse function defined in Eq. (3).
The dispersive nature of the plasma is encoded in the refractive index,
$n^{2}=1-\frac{\omega_{p}^{2}}{\omega^{2}},$ (35)
where $\omega$ denotes the photon frequency measured by a local static observer. The plasma frequency itself is determined by the electron number density $N(r)$ through
$\omega_{p}^{2}(r)=\frac{4\pi e^{2}N(r)}{m_{e}},$ (36)
with $e$ and $m_{e}$ representing the electron charge and mass, respectively. A photon detected at spatial infinity with frequency $\omega_{0}$ undergoes gravitational redshift as it propagates inward, such that
$\omega(r)=\frac{\omega_{0}}{\sqrt{f(r)}}.$ (37)
The vacuum limit is recovered smoothly by setting $\omega_{p}\to 0$, in which case the Hamiltonian reduces to the standard null geodesic condition $g^{\alpha\beta}p_{\alpha}p_{\beta}=0$.
Restricting attention to the equatorial plane ($\theta=\pi/2$), Hamilton’s equations yield the equations of motion
$\dot{t}=-\frac{p_{t}}{f(r)},\qquad\dot{r}=f(r)p_{r},\qquad\dot{\varphi}=\frac{p_{\varphi}}{r^{2}},$ (38)
where the overdot denotes differentiation with respect to an affine parameter. Imposing the Hamiltonian constraint leads to the radial orbit equation
$\frac{dr}{d\varphi}=r\sqrt{f(r)}\sqrt{h^{2}(r)\frac{\omega_{0}^{2}}{p_{\varphi}^{2}}-1},$ (39)
where the optical impact function $h(r)$ is defined as Sucu and Sakallı (2025b)
$h^{2}(r)=r^{2}\left(\frac{1}{f(r)}-\frac{\omega_{p}^{2}(r)}{\omega_{0}^{2}}\right).$ (40)
Unstable circular photon orbits, which determine the photon sphere and the boundary of the black hole shadow, correspond to extrema of the impact function $h^{2}(r)$. The photon-sphere radius $r_{p}$ is therefore obtained from the condition
$\frac{d}{dr}\left[r^{2}\left(\frac{1}{f(r)}-\frac{\omega_{p}^{2}(r)}{\omega_{0}^{2}}\right)\right]_{r=r_{p}}=0,$ (41)
or equivalently,
$\frac{2}{f(r_{p})}-\frac{2\omega_{p}^{2}(r_{p})}{\omega_{0}^{2}}-r_{p}\left(\frac{f^{\prime}(r_{p})}{f^{2}(r_{p})}+\frac{2\omega_{p}(r_{p})\omega_{p}^{\prime}(r_{p})}{\omega_{0}^{2}}\right)=0.$ (42)
This condition generalizes the standard vacuum photon-sphere equation by incorporating plasma effects through both the local plasma frequency $\omega_{p}(r_{p})$ and its radial gradient $\omega_{p}^{\prime}(r_{p})$. As a result, the photon-sphere radius becomes explicitly frequency dependent and sensitive to the plasma distribution, providing a direct channel through which environmental effects can modify observable shadow and lensing properties.
VI.2 Homogeneous versus inhomogeneous plasma distributions
The influence of plasma on photon-sphere properties depends crucially on the spatial structure of the medium. To disentangle intrinsic spacetime effects from environmental contributions, we examine two representative scenarios: a homogeneous plasma with constant plasma frequency and an inhomogeneous plasma with a radially stratified density profile.
For a spatially uniform plasma, the plasma frequency is constant, $\omega_{p}=\text{const.}$, and the gradient term in Eq. (42) vanishes. The photon-sphere condition then reduces to
$\frac{2}{f(r_{p})}-\frac{2\omega_{p}^{2}}{\omega_{0}^{2}}=r_{p}\frac{f^{\prime}(r_{p})}{f^{2}(r_{p})},$ (43)
which can be solved numerically for the Kalb-Ramond-ModMax metric function (3) at fixed values of the plasma-to-photon frequency ratio $\omega_{p}^{2}/\omega_{0}^{2}$.
Table 2: Photon sphere radius $r_{p}$ for KR-ModMax black holes in homogeneous plasma with fixed $M=1$, $\ell=0.2$, and $Q=0.5$. The ordinary branch ($\zeta=+1$) yields smaller photon sphere radii due to the repulsive electromagnetic contribution, while the phantom branch ($\zeta=-1$) produces larger $r_{p}$ from enhanced gravitational focusing. Increasing plasma density ($\omega_{p}^{2}/\omega_{0}^{2}$) uniformly shifts $r_{p}$ outward for both branches.
$\zeta$ $\gamma$ $\omega_{p}^{2}/\omega_{0}^{2}$ $r_{p}$
Ordinary Branch ($\zeta=+1$)
+1+1 0 $0.00$ $2.103$
+1+1 0 $0.05$ $2.118$
+1+1 0 $0.10$ $2.134$
+1+1 0 $0.20$ $2.171$
+1+1 11 $0.00$ $2.300$
+1+1 11 $0.05$ $2.317$
+1+1 11 $0.10$ $2.334$
+1+1 11 $0.20$ $2.374$
Phantom Branch ($\zeta=-1$)
1-1 0 $0.00$ $2.637$
1-1 0 $0.05$ $2.656$
1-1 0 $0.10$ $2.677$
1-1 0 $0.20$ $2.723$
1-1 11 $0.00$ $2.492$
1-1 11 $0.05$ $2.510$
1-1 11 $0.10$ $2.530$
1-1 11 $0.20$ $2.573$
The results summarized in Table 2 reveal several robust trends. First, the presence of plasma systematically shifts the photon sphere outward for both branches. As the ratio $\omega_{p}^{2}/\omega_{0}^{2}$ increases, the effective propagation speed of photons decreases, weakening curvature-induced focusing and pushing unstable circular orbits to larger radii. Quantitatively, increasing $\omega_{p}^{2}/\omega_{0}^{2}$ from 0 to $0.20$ leads to a shift $\Delta r_{p}\approx 0.07$ for the ordinary branch and $\Delta r_{p}\approx 0.09$ for the phantom branch.
Second, the branch dichotomy remains clearly visible in homogeneous plasma. For identical values of $(\ell,\gamma)$, the phantom branch ($\zeta=-1$) consistently produces larger photon-sphere radii than the ordinary branch ($\zeta=+1$). In vacuum and at $\gamma=0$, this difference reaches $\Delta r_{p}\approx 0.53$, corresponding to an enhancement of roughly $25\%$. Importantly, this separation persists across all plasma densities, indicating that it originates from the intrinsic sign reversal of the electromagnetic sector rather than from plasma effects.
Finally, the ModMax parameter $\gamma$ affects the two branches in qualitatively different ways. In the ordinary branch, increasing $\gamma$ suppresses the effective charge contribution through the factor $e^{-\gamma}$ and shifts the geometry toward the Schwarzschild-Kalb-Ramond limit, resulting in a moderate increase of $r_{p}$. In the phantom branch, the same increase in $\gamma$ partially suppresses the enhanced focusing induced by the phantom charge, leading instead to a decrease of the photon-sphere radius. This opposite response provides an additional handle for distinguishing between the two branches in weakly stratified environments.
Now, astrophysical plasmas are typically radially stratified rather than homogeneous. To model this situation, we consider a power-law plasma profile,
$\omega_{p}^{2}(r)=\frac{\kappa_{0}}{r^{\alpha}},$ (44)
with $\alpha=1$, which captures the characteristic $1/r$ falloff expected in stellar winds and accretion flows. Substituting this profile into Eq. (42) yields
$\frac{2}{f(r_{p})}-\frac{2\kappa_{0}}{r_{p}\omega_{0}^{2}}-r_{p}\left(\frac{f^{\prime}(r_{p})}{f^{2}(r_{p})}+\frac{2\kappa_{0}^{2}}{r_{p}^{3}\omega_{0}^{2}}\right)=0,$ (45)
which must again be solved numerically.
Table 3: Photon sphere radius $r_{p}$ for KR-ModMax black holes in inhomogeneous plasma ($\alpha=1$) with fixed $M=1$, $Q=0.5$, $\kappa_{0}=1$, and $\omega_{0}=0.5$. The KR parameter $\ell$ dominates the photon sphere structure, producing a four-fold increase in $r_{p}$ as $\ell$ increases from $0.1$ to $0.8$. The ModMax branch parameter $\zeta$ and nonlinearity $\gamma$ produce only marginal corrections at the third decimal place.
$\ell$ $\zeta$ $\gamma$ $r_{p}$
Weak KR Deformation ($\ell=0.1$)
$0.1$ +1+1 0 $4.734$
$0.1$ +1+1 11 $4.773$
$0.1$ 1-1 0 $4.860$
$0.1$ 1-1 11 $4.819$
Moderate KR Deformation ($\ell=0.5$)
$0.5$ +1+1 0 $8.447$
$0.5$ +1+1 11 $8.452$
$0.5$ 1-1 0 $8.465$
$0.5$ 1-1 11 $8.459$
Strong KR Deformation ($\ell=0.8$)
$0.8$ +1+1 0 $20.762$
$0.8$ +1+1 11 $20.763$
$0.8$ 1-1 0 $20.765$
$0.8$ 1-1 11 $20.764$
The results collected in Table 3 display a markedly different hierarchy of effects. In this case, the photon-sphere radius is dominated by the Kalb-Ramond parameter $\ell$, with only marginal sensitivity to the ModMax sector. As $\ell$ increases from $0.1$ to $0.8$, the photon-sphere radius grows by more than a factor of four, reflecting the strong amplification induced by the asymptotic rescaling $(1-\ell)^{-1}$ in combination with the radially varying refractive index.
On the other hand, variations in the branch parameter $\zeta$ and in the ModMax parameter $\gamma$ produce only minute shifts in $r_{p}$, typically at the level $\Delta r_{p}\sim 10^{-3}$-$10^{-2}$. For instance, at $\ell=0.8$ the difference between the ordinary and phantom branches is reduced to $\Delta r_{p}\approx 0.003$, compared to $\Delta r_{p}\approx 0.53$ in the homogeneous case. This strong suppression indicates that the plasma gradient dominates the effective optical potential, effectively masking the sign-sensitive electromagnetic contributions.
We close this subsection by discussing the physical and observational implications. The contrast between homogeneous and inhomogeneous plasma regimes has direct implications for black-hole imaging. In homogeneous or weakly stratified media, the photon-sphere radius retains clear sensitivity to the electromagnetic branch structure, and the sizable separation between ordinary and phantom configurations could, in principle, be probed through precise measurements of shadow sizes or frequency-dependent lensing. In strongly stratified environments, however, plasma effects and Kalb-Ramond-induced rescaling dominate photon dynamics, rendering branch discrimination observationally challenging.
At the same time, the pronounced sensitivity of the photon-sphere radius to the Lorentz-symmetry-breaking parameter $\ell$ in inhomogeneous plasma suggests that high-resolution, multi-frequency observations could place meaningful constraints on such deformations, provided that plasma properties are modeled accurately. Overall, these results reveal the necessity of incorporating realistic plasma distributions when attempting to extract fundamental gravitational or electromagnetic physics from horizon-scale observations.
VII Tidal forces and branch-dependent deformations
Tidal effects acting on extended bodies provide a direct probe of spacetime curvature beyond the idealized motion of pointlike test particles. While geodesic trajectories encode the global structure of a gravitational field, finite-sized objects respond to local curvature gradients through differential accelerations governed by the Riemann tensor. In strong-field regimes, these tidal deformations can exhibit qualitative features that are invisible at the level of geodesic motion alone.
In the Kalb-Ramond-ModMax framework, the combined presence of Lorentz-symmetry-breaking effects and nonlinear electrodynamics modifies the radial dependence of the metric function relative to standard electrovacuum solutions. As a result, the tidal response of infalling matter can depart significantly from the familiar Schwarzschild or Reissner-Nordström behavior. In this section, we analyze tidal forces in the ordinary and phantom branches and identify branch-dependent deformation patterns with no counterpart in general relativity.
VII.1 Geodesic deviation and tidal forces
The relative acceleration between neighboring freely falling worldlines is governed by the geodesic deviation equation Cordeiro et al. (2025)
$\frac{D^{2}\eta^{\mu}}{D\tau^{2}}=-R^{\mu}{}_{\nu\alpha\beta}\,u^{\nu}u^{\alpha}\eta^{\beta},$ (46)
where $\eta^{\mu}$ denotes the separation vector between nearby geodesics, $u^{\mu}$ is the four-velocity of the reference worldline, and $\tau$ is the proper time. The Riemann curvature tensor acts as the mediator of tidal interactions, converting spacetime geometry into physically measurable distortions of extended bodies.
To obtain physically transparent tidal forces, it is convenient to project the curvature tensor onto an orthonormal tetrad adapted to static observers. Introducing the basis Chandrasekhar (1985); Poisson (2004)
$e_{\hat{0}}$ $=f^{-1/2}\,\partial_{t},$
$e_{\hat{1}}$ $=f^{1/2}\,\partial_{r},$
$e_{\hat{2}}$ $=\frac{1}{r}\,\partial_{\theta},$
$e_{\hat{3}}$ $=\frac{1}{r\sin\theta}\,\partial_{\varphi},$ (47)
the metric locally reduces to the Minkowski form $\eta_{\hat{a}\hat{b}}=\mathrm{diag}(-1,1,1,1)$.
In this frame, tidal forces are encoded in the electric part of the Riemann tensor,
$\mathcal{E}_{\hat{i}\hat{j}}=R_{\hat{0}\hat{i}\hat{0}\hat{j}}$, (48)
which directly determines the relative acceleration of nearby freely falling particles. For any static, spherically symmetric spacetime, the independent nonvanishing components reduce to
$\mathcal{E}_{\hat{1}\hat{1}}=-\frac{1}{2}f^{\prime\prime}(r),\qquad\mathcal{E}_{\hat{2}\hat{2}}=\mathcal{E}_{\hat{3}\hat{3}}=-\frac{f^{\prime}(r)}{2r},$ (49)
with spherical symmetry enforcing equality of the two angular components.
For the Kalb-Ramond-ModMax black hole, the first and second radial derivatives of the metric function (3) are given by
$f^{\prime}(r)$ $=\frac{2M}{r^{2}}-\frac{2\zeta Q^{2}e^{-\gamma}}{(1-\ell)^{2}r^{3}},$ (50)
$f^{\prime\prime}(r)$ $=-\frac{4M}{r^{3}}+\frac{6\zeta Q^{2}e^{-\gamma}}{(1-\ell)^{2}r^{4}}.$ (51)
Substituting these expressions into the geodesic deviation equation yields the explicit tidal accelerations,
$\frac{D^{2}\eta_{\hat{1}}}{D\tau^{2}}$ $=\left(\frac{2M}{r^{3}}-\frac{3\zeta Q^{2}e^{-\gamma}}{(1-\ell)^{2}r^{4}}\right)\eta_{\hat{1}}$, (52)
$\frac{D^{2}\eta_{\hat{\imath}}}{D\tau^{2}}$ $=\left(-\frac{M}{r^{3}}+\frac{\zeta Q^{2}e^{-\gamma}}{(1-\ell)^{2}r^{4}}\right)\eta_{\hat{\imath}},\qquad\hat{\imath}=\hat{2},\hat{3}$. (53)
At large radii, the leading $r^{-3}$ terms dominate, recovering the familiar Schwarzschild tidal pattern characterized by radial stretching accompanied by angular compression. Closer to the black hole, however, the charge-dependent terms scale more steeply with radius and can significantly reshape the tidal profile. The manner in which these corrections compete or reinforce the Schwarzschild contribution depends crucially on the electromagnetic branch parameter $\zeta$, setting the stage for qualitatively distinct tidal behaviors in the ordinary and phantom sectors.
VII.2 Tidal inversion and physical implications
A distinctive feature of the Kalb-Ramond-ModMax black hole is the possibility of tidal-force inversion, namely the vanishing and subsequent sign reversal of tidal accelerations at finite radii outside the event horizon. This phenomenon has no analogue in the Schwarzschild spacetime and arises from the competition between the mass-induced curvature and the charge-dependent contributions controlled by the Kalb-Ramond and ModMax sectors.
Figure 9: Radial tidal acceleration $T_{\mathrm{rad}}$ as a function of $r$ for different values of the Lorentz-violating parameter $\ell$ in the ordinary branch ($\zeta=+1$). Increasing $\ell$ shifts the zero-crossing radius $R_{\mathrm{rad}}$ to larger values and enhances the tidal forces at intermediate radii. Parameters: $M=1$, $Q=0.8$, $\gamma=0.5$.
The sensitivity of tidal forces to the Lorentz-symmetry-breaking parameter $\ell$ is illustrated in Fig. 9 for the ordinary branch. As $\ell$ increases, the factor $(1-\ell)^{-2}$ amplifies the charge-dependent terms in Eqs. (52)-(53) relative to the Schwarzschild contribution. This produces two correlated effects: the magnitude of the tidal acceleration at fixed intermediate radii increases, and the radial location at which the radial tidal force vanishes is shifted outward. Physically, stronger Lorentz violation enlarges the region where charge-induced corrections dominate the tidal response, thereby extending the domain of tidal inversion.
Figure 10: Tidal balance radii $R_{\mathrm{rad}}$ (blue) and $R_{\mathrm{ang}}$ (red) as functions of the electric charge $Q$ for the ordinary ModMax branch ($\zeta=+1$). Both radii scale quadratically with $Q$ and preserve the universal ratio $R_{\mathrm{rad}}/R_{\mathrm{ang}}=3/2$, independent of model parameters. Parameters: $M=1$, $\ell=0.1$, $\gamma=0.5$.
Table 4: Tidal balance radii for the KR-ModMax BH in the ordinary branch ($\zeta=+1$), with $M=1$ and $Q=0.8$. The ratio $R_{\mathrm{rad}}/R_{\mathrm{ang}}=3/2$ holds universally.
$\ell$ $\gamma$ $R_{\mathrm{rad}}$ $R_{\mathrm{ang}}$
0.00 0.0 0.9600 0.6400
0.00 0.5 0.5823 0.3882
0.00 1.0 0.3532 0.2354
0.00 1.5 0.2142 0.1428
0.10 0.0 1.1852 0.7901
0.10 0.5 0.7189 0.4792
0.10 1.0 0.4360 0.2907
0.10 1.5 0.2645 0.1763
0.20 0.0 1.5000 1.0000
0.20 0.5 0.9098 0.6065
0.20 1.0 0.5518 0.3679
0.20 1.5 0.3347 0.2231
0.30 0.0 1.9592 1.3061
0.30 0.5 1.1883 0.7922
0.30 1.0 0.7207 0.4805
0.30 1.5 0.4372 0.2914
The characteristic transition scales are quantified by the tidal balance radii $R_{\mathrm{rad}}$ and $R_{\mathrm{ang}}$, defined by the vanishing of the radial and angular tidal components, respectively. Their dependence on the electric charge and model parameters is displayed in Fig. 10 and summarized numerically in Table 4. The data reveal systematic trends: both balance radii increase with $\ell$ due to the universal $(1-\ell)^{-2}$ enhancement, while increasing the ModMax parameter $\gamma$ suppresses the charge contribution through the factor $e^{-\gamma}$, shifting the inversion radii inward. For representative parameter choices, the balance radii span the range $R_{\mathrm{rad}},R_{\mathrm{ang}}\sim 0.1$-22 in units of the black-hole mass, indicating that tidal inversion may occur either close to the horizon or well outside it, depending on the regime.
The existence and physical relevance of these inversion scales depend crucially on the electromagnetic branch parameter $\zeta$. For the ordinary ModMax branch ($\zeta=+1$), the charge contribution enters the tidal tensor with a sign opposite to the Schwarzschild term at sufficiently small radii. As shown in Fig. 11, radial tidal stretching weakens as the black hole is approached, vanishes at $R_{\mathrm{rad}}$, and is subsequently replaced by radial compression. Simultaneously, angular compression is softened and reverses at $R_{\mathrm{ang}}$. This regulated tidal environment represents a qualitative departure from the classical spaghettification scenario and suggests that the Kalb-Ramond-ModMax corrections act as an effective curvature-softening mechanism. Such behavior could influence the coherence, deformation, and survivability of extended structures—such as accretion streams or plasma filaments—in the strong-field region.
Figure 11: Radial (solid lines) and angular (dashed lines) tidal acceleration components for KR-ModMax black holes, evaluated in an orthonormal frame. Blue curves correspond to the ordinary branch ($\zeta=+1$), while red curves represent the phantom branch ($\zeta=-1$). The horizontal gray dashed line marks vanishing tidal acceleration. For the ordinary branch, both components cross zero at finite radii, whereas no zero crossing occurs for the phantom branch. Parameters: $M=1$, $Q=0.8$, $\ell=0.1$, $\gamma=0.5$.
On the other hand, for the phantom branch ($\zeta=-1$) the electromagnetic contribution reinforces the Schwarzschild tidal field rather than opposing it. In this case, both $R_{\mathrm{rad}}$ and $R_{\mathrm{ang}}$ become negative and therefore lose any direct geometric meaning. Radial stretching and angular compression persist throughout the exterior region and grow monotonically toward the horizon, producing tidal stresses that are even stronger than those encountered in standard general relativity. This behavior reflects the exotic energy content of the phantom ModMax sector and indicates a dynamically harsher environment for infalling extended matter.
The aforementioned dichotomy between the two branches reveals the sensitivity of tidal observables to the underlying nonlinear electrodynamic structure. While both branches are mathematically admissible solutions of the Einstein-Kalb-Ramond-ModMax equations, tidal dynamics provide a physically transparent discriminator between them. When combined with thermodynamic behavior and optical signatures such as gravitational lensing and photon-sphere properties, tidal effects furnish an independent and complementary probe for assessing the physical viability of ordinary versus phantom regimes in this class of Lorentz-violating black-hole models.
From a physical standpoint, the existence or absence of tidal inversion constitutes an independent diagnostic of branch viability. In the ordinary ModMax branch, the partial cancellation between the Schwarzschild and electromagnetic contributions leads to a regulated strong-field tidal environment, in which extreme spaghettification is softened within a finite radial interval outside the horizon. By contrast, the phantom branch exhibits no such regulation: tidal stretching and angular compression are uniformly enhanced, producing a dynamically harsher environment for extended matter.
Importantly, this distinction is entirely local and does not rely on thermodynamic equilibrium or asymptotic optical measurements. When combined with the thermodynamic phase structure and the branch-dependent lensing and photon-sphere properties discussed in previous sections, tidal forces therefore provide a complementary and physically transparent probe of the underlying nonlinear electrodynamic sector.
VIII Conclusions
In this work we have explored a charged black-hole solution arising from the simultaneous presence of a Kalb-Ramond (KR) three-form condensate and ModMax nonlinear electrodynamics, focusing on the physical implications of the resulting branch structure. The motivation behind this study was twofold. On one hand, both ingredients are well motivated independently: the Kalb-Ramond field naturally appears in string-inspired effective theories and induces controlled Lorentz-symmetry-breaking effects, while ModMax electrodynamics represents the unique conformal and duality-invariant deformation of Maxwell theory. On the other hand, their combined impact on black-hole physics has remained largely unexplored, despite the fact that each sector modifies gravity in a distinct and potentially observable way. Our aim was therefore to assess whether their interplay leads merely to quantitative corrections, or instead to qualitatively new physical behavior.
The resulting KR-ModMax black hole is governed, beyond mass and charge, by three deformation parameters: the Lorentz-symmetry-breaking parameter $\ell$, the ModMax nonlinearity parameter $\gamma$, and the discrete branch selector $\zeta=\pm 1$. A central outcome of our analysis is that the branch parameter $\zeta$ acts as a genuine physical discriminator. Already at the level of the spacetime geometry, the ordinary branch ($\zeta=+1$) admits the familiar hierarchy of non-extremal, extremal, and naked configurations, while the phantom branch ($\zeta=-1$) generically supports a single-horizon structure. This geometric distinction propagates coherently through all physical sectors that we investigated, providing a unifying theme for the entire analysis.
We examined the thermodynamic properties of the KR-ModMax black hole within the non-extensive Tsallis entropy framework, which is particularly suited to gravitational systems characterized by long-range correlations. The Hawking temperature exhibits branch-dependent behavior: charge effects suppress the temperature in the ordinary branch, allowing for extremal configurations, whereas in the phantom branch the temperature is uniformly enhanced and never vanishes. The Tsallis deformation parameter $\delta$ reshapes the thermodynamic landscape, modifying the internal energy, Helmholtz free energy, and effective pressure. The heat capacity proved especially informative: while the ordinary branch can approach regimes suggestive of thermodynamic stabilization, the phantom branch remains dominated by negative heat capacity, indicating persistent local instability. These features are further reinforced by the Joule-Thomson expansion, where well-defined inversion curves appear primarily in the ordinary branch, whereas the phantom sector exhibits a more restricted and less structured cooling-heating behavior.
Optical properties provide an independent and complementary window into the branch structure. Using the Gauss-Bonnet approach to weak gravitational lensing, we found that the ordinary branch generally produces reduced deflection angles compared to the Schwarzschild case, reflecting the partially repulsive role of the electromagnetic sector. In contrast, the phantom branch yields systematically enhanced deflection, since the sign-reversed electromagnetic contribution reinforces gravitational focusing. This distinction persists, in a modified form, when photon propagation is studied in plasma environments. In homogeneous plasma, branch-dependent shifts of the photon-sphere radius remain visible, while in strongly stratified plasma distributions the optical structure becomes dominated by the Lorentz-symmetry-breaking parameter $\ell$, with ModMax effects largely suppressed. This hierarchy reveals the importance of accurate plasma modeling when attempting to extract fundamental physics from black-hole shadow observations.
Tidal forces offer a direct probe of the local curvature structure that is independent of both thermodynamic equilibrium and asymptotic optical measurements. By analyzing geodesic deviation in an orthonormal frame, we identified two characteristic tidal balance radii at which radial and angular tidal components vanish. For the ordinary branch, these radii are positive and physically meaningful, signaling a partial inversion of the classical spaghettification pattern: radial stretching and angular compression are softened and can reverse inside a finite region outside the horizon. Remarkably, the ratio of these radii is universal and independent of all model parameters. The phantom branch, by contrast, admits no such tidal inversion; tidal forces are uniformly intensified as the horizon is approached. This sharp dichotomy provides a physically transparent criterion for assessing branch viability and reinforces the conclusions drawn from thermodynamic and optical analyses.
Taken together, our results demonstrate that the KR-ModMax black hole is not merely a perturbative deformation of Reissner-Nordström geometry, but a theoretically motivated system with a rich and internally consistent physical structure. The ordinary branch retains continuity with standard charged black holes while introducing controlled, potentially observable deviations. The phantom branch, although mathematically admissible, displays a combination of enhanced lensing, stronger tidal forces, and persistent thermodynamic instability that may challenge its physical realizability, or alternatively render it particularly easy to distinguish observationally.
Extending the analysis to rotating configurations would greatly enhance the astrophysical relevance of the model. Moreover, quasinormal modes and gravitational-wave ringdown signatures could provide additional observational discriminants. Additionally, a systematic comparison between Tsallis entropy and other generalized entropy formalisms may further illuminate the role of non-extensivity in black-hole thermodynamics. Finally, embedding the KR-ModMax framework within a more fundamental string-theoretic or supergravity setting could clarify the microscopic origin and allowed ranges of the parameters $\ell$, $\gamma$, and $\zeta$. All these studies will be performed in future projects.
In conclusion, the combined presence of Kalb-Ramond Lorentz-symmetry breaking and ModMax nonlinear electrodynamics leads to a black-hole phenomenology that is both rich and sharply structured. The branch parameter $\zeta$ emerges as the key organizing principle, controlling geometry, thermodynamics, optics, and tidal dynamics in a unified way, and offering multiple, complementary pathways for confronting this class of models with observations.
Acknowledgements.
İ. S. and E. S. thank TÜBİTAK, ANKOS, and SCOAP3 for academic support. The authors acknowledge the contribution of the LISA Cosmology Working Group (CosWG), as well as support from the COST Actions CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse) - CA23130, Bridging high and low energies in search of quantum gravity (BridgeQG) and CA21106 - COSMIC WISPers in the Dark Universe: Theory, astrophysics and experiments (CosmicWISPers).
Appendix A Equatorial embeddings and spatial geometry
(a) $\zeta=-1$, $\ell=0.2$, $\gamma=0.5$, $Q=0.2$ ($r_{+}=1.62$). Weak Lorentz-symmetry-breaking and ModMax effects produce a wide, shallow embedding surface with a gradual approach to the horizon.
(b) $\zeta=-1$, $\ell=0.5$, $\gamma=3.0$, $Q=0.5$ ($r_{+}=1.02$). Strong ModMax suppression significantly reduces the effective charge, yielding a compact horizon and a steep near-horizon geometry.
(c) $\zeta=-1$, $\ell=0.7$, $\gamma=1.0$, $Q=0.8$ ($r_{+}=1.24$). Large Kalb-Ramond deformation dominates the geometry, producing pronounced curvature and enhanced asymptotic normalization.
(d) $\zeta=-1$, $\ell=0.4$, $\gamma=2.0$, $Q=0.1$ ($r_{+}=1.20$). Small-charge configuration approaching the uncharged KR-Schwarzschild limit, with the single-horizon topology preserved.
Figure 12: Three-dimensional isometric embedding diagrams of the equatorial plane ($\theta=\pi/2$, $t=\mathrm{const.}$) for phantom-branch ($\zeta=-1$) KR-ModMax black holes. Black rings mark the event-horizon radius $r=r_{+}$, while blue spirals represent timelike infalling trajectories. The depth and curvature of the embedding surface provide a geometric visualization of the combined effects of Lorentz symmetry breaking and nonlinear electrodynamics.
To provide a geometric visualization of the spatial curvature induced by the Kalb-Ramond-ModMax fields, we construct isometric embeddings of the equatorial plane ($\theta=\pi/2$, $t=\text{const.}$) into three-dimensional Euclidean space. Although such embeddings do not carry direct observational significance, they offer an intuitive representation of how the underlying spacetime geometry is deformed by the combined effects of Lorentz symmetry breaking and nonlinear electrodynamics.
The induced two-dimensional metric on the equatorial slice is given by
$d\sigma^{2}=\frac{dr^{2}}{f(r)}+r^{2}\,d\phi^{2},$ (54)
where $f(r)$ denotes the lapse function of the KR-ModMax black hole. This metric can be embedded in Euclidean space using the standard cylindrical parametrization $(r,\phi)\mapsto(r\cos\phi,r\sin\phi,z(r))$, where the embedding function $z(r)$ satisfies
$\left(\frac{dz}{dr}\right)^{2}=\frac{1}{f(r)}-1.$ (55)
Real solutions exist in regions where $f(r)<1$, leading to the familiar funnel-shaped surfaces that encode the effective gravitational potential of the black hole.
In Fig. 12 we present representative equatorial embedding diagrams for four configurations belonging to the phantom ModMax branch. The surfaces extend outward from the event horizon radius $r_{+}$ (indicated by a black ring) and illustrate how spatial curvature accumulates as the horizon is approached. The vertical direction measures the embedding height $z(r)$, with steeper funnels corresponding to stronger curvature gradients. Superimposed blue spiral curves depict timelike geodesics of infalling test particles, included to visualize the qualitative motion of matter toward the horizon.
Several qualitative features can be identified. In Panel (a) ($\ell=0.2$, $\gamma=0.5$, $Q=0.2$, $r_{+}=1.62$) we see a modest Kalb-Ramond deformation produces a wide, gently sloping funnel, indicative of a relatively weak curvature enhancement. In Panel (b) ($\ell=0.5$, $\gamma=3.0$, $Q=0.5$, $r_{+}=1.02$) we observe a strong ModMax suppression ($e^{-\gamma}\ll 1$) yields a compact horizon and a sharply curved near-horizon geometry. In Panel (c) ($\ell=0.7$, $\gamma=1.0$, $Q=0.8$, $r_{+}=1.24$) we can see a large Lorentz-symmetry-breaking parameter amplifies asymptotic deviations, producing pronounced geometric flattening away from the horizon. Finally, in Panel (d) ($\ell=0.4$, $\gamma=2.0$, $Q=0.1$, $r_{+}=1.20$) we have a small electric charge with intermediate parameters illustrates the smooth transition toward the uncharged KR-Schwarzschild limit.
An important structural aspect revealed by these embeddings is the single-horizon topology of the phantom branch. In contrast to Reissner-Nordström-type geometries, no inner Cauchy horizon is present. As a result, the spacetime avoids the mass-inflation instability that typically afflicts charged black holes at the inner horizon. While a full dynamical stability analysis lies beyond the scope of this work, the absence of an inner horizon suggests that phantom KR-ModMax black holes may exhibit improved robustness against perturbations in the deep interior.

### v2

@@
 Black-hole thermodynamics is analyzed within the Tsallis non-extensive framework, revealing branch-dependent stability and Joule-Thomson behavior.
-Weak gravitational lensing, photon propagation in plasma, and tidal forces are then studied, revealing clear optical and strong-field signatures that distinguish the two branches.
-In particular, the ordinary branch exhibits finite-radius tidal inversion, absent in the phantom sector.
-Our results demonstrate that the combined Kalb-Ramond and ModMax effects lead to a rich and observationally distinguishable black-hole phenomenology.
+Weak gravitational lensing is analyzed using the Ono-Ishihara-Asada extension of the Gauss-Bonnet theorem, which is required by the non-Euclidean asymptotic structure of the Kalb-Ramond optical geometry and yields a negative topological correction that reduces light bending relative to the Schwarzschild baseline.
+Photon propagation in plasma and tidal forces are also studied, revealing clear optical and strong-field signatures that distinguish the two branches.
 I Introduction
@@
 (2024).
+However, the standard formulation assumes asymptotically Euclidean optical geometry, a condition that is violated for the KR-ModMax metric due to the modified lapse at infinity.
+In such cases, the Ono-Ishihara-Asada (OIA) generalisation Ono et al.
+(2017); Ishihara et al.
+(2016); Jusufi et al.
+(2017) must be employed, which introduces a topological boundary correction encoding the non-flat asymptotic structure.
 Moreover, realistic astrophysical environments are typically permeated by plasma, which modifies photon propagation through dispersive effects.
@@
 Although both the Kalb-Ramond condensate and the ModMax nonlinearity primarily affect the near-horizon geometry, their influence propagates to large distances through the optical structure of spacetime.
-To capture these effects in a coordinate-independent manner, we employ the Gauss-Bonnet (GB) theorem approach Gibbons and Werner (2008); Sucu and Sakalli (2025c); Sucu and Sakallı (2025c), which interprets the deflection of light as a global geometric effect governed by the integrated Gaussian curvature of an associated optical manifold.
-V.1 Optical geometry and Gauss-Bonnet method
-In static and spherically symmetric spacetimes, null geodesics can be equivalently described as spatial geodesics on an effective two-dimensional Riemannian manifold known as the optical geometry ABRAMOWICZ et al.
-(1988); Perlick (2006).
-Restricting the Kalb-Ramond-ModMax metric (1) to the equatorial plane ($\theta=\pi/2$), the corresponding optical line element takes the form
+A key feature of the metric (3) is its non-standard asymptotic behaviour, $f(r)\to 1/(1-\ell)\neq 1$ as $r\to\infty$, which renders the standard Gibbons-Werner formulation of the Gauss-Bonnet theorem inapplicable Gibbons and Werner (2008); Jusufi et al.
+(2017); Ishihara et al.
+(2016).
+Instead, the deflection angle must be computed using the Ono-Ishihara-Asada (OIA) extension Ono et al.
+(2017); Ishihara et al.
+(2016), which incorporates a topological boundary correction arising from the non-Euclidean asymptotic structure of the optical manifold.
+V.1 Non-asymptotic flatness and coordinate transformation
+The optical metric associated with a static, spherically symmetric spacetime (1) restricted to the equatorial plane ($\theta=\pi/2$) takes the form
 $d\sigma^{2}=\frac{dr^{2}}{f^{2}(r)}+\frac{r^{2}}{f(r)}\,d\phi^{2},$ (25)
@@
 (2015).
-The intrinsic curvature of the optical manifold is characterized by the Gaussian curvature $K$, which in two dimensions is fully determined by a single independent component of the Riemann tensor Sucu and Sakallı (2025a), namely
-$K=\frac{R_{r\phi r\phi}}{\det(g_{\rm opt})}.$ (26)
-In the weak-field region ($r\gg r_{h}$), we expand the metric function asymptotically and retain terms up to $\mathcal{O}(r^{-5})$. This yields the following expression for the Gaussian curvature:
-$\!\!\!\!\!K\simeq\frac{3\zeta Q^{2}}{r^{4}(1-\ell)^{4}e^{\gamma}}-\frac{3\zeta Q^{2}\ell}{r^{4}(1-\ell)^{4}e^{\gamma}}+\frac{2M\ell^{3}}{r^{3}(1-\ell)^{4}}$
-$\ \ \ \ -\frac{6M\ell^{2}}{r^{3}(1-\ell)^{4}}+\frac{6M\ell}{r^{3}(1-\ell)^{4}}-\frac{2M}{r^{3}(1-\ell)^{4}}$
-$\ \ \ \ +\frac{3M^{2}\ell^{4}}{r^{4}(1-\ell)^{4}}-\frac{12M^{2}\ell^{3}}{r^{4}(1-\ell)^{4}}+\frac{18M^{2}\ell^{2}}{r^{4}(1-\ell)^{4}}$
-$\ \ \ \ -\frac{12M^{2}\ell}{r^{4}(1-\ell)^{4}}+\frac{3M^{2}}{r^{4}(1-\ell)^{4}}.$ (27)
-The leading $r^{-3}$ contributions originate from the mass sector, while the $r^{-4}$ terms encode both second-order mass corrections and the charge-dependent ModMax contributions, controlled by the branch parameter $\zeta$ and the suppression factor $e^{-\gamma}$.
-The Kalb-Ramond parameter $\ell$ enters universally through the prefactor $(1-\ell)^{-4}$, amplifying all curvature contributions as $\ell$ approaches unity.
-The Gauss-Bonnet theorem provides a direct link between this intrinsic curvature and the total deflection angle experienced by a light ray.
-Applied to a non-singular domain $\mathcal{D}_{\infty}$ exterior to the photon trajectory, the theorem yields the coordinate-independent expression Gibbons and Werner (2008)
-$\beta=-\iint_{\mathcal{D}_{\infty}}K\,dS,$ (28)
-where $dS$ denotes the surface element of the optical manifold.
-In the weak-lensing approximation, the photon path is well approximated by the unperturbed straight-line trajectory $r=b/\sin\phi$, allowing the integral to be evaluated analytically.
-The resulting deflection angle, together with its branch-dependent structure and limiting behavior, is discussed in the following subsection.
-V.2 Weak-field deflection angle and branch-dependent signatures
-Within the Gauss-Bonnet framework, the total deflection angle experienced by a light ray with impact parameter $b$ is determined by the surface integral of the Gaussian curvature over the exterior optical domain $\mathcal{D}_{\infty}$, namely
-$\beta=-\iint_{\mathcal{D}_{\infty}}K\,dS,$ (29)
-where the integration region is bounded by the photon trajectory and an asymptotic reference curve Gibbons and Werner (2008).
-In the weak-field regime, the photon path is well approximated by the unperturbed straight-line trajectory $r=b/\sin\phi$, which allows the integral to be evaluated analytically.
-Carrying out the integration using the asymptotic expansion of the Gaussian curvature given in Eq. (27), we obtain the vacuum deflection angle for the Kalb-Ramond-ModMax black hole,
-$\!\!\!\!\!\!\beta_{\rm vac}\approx\frac{4M}{b(1-\ell)^{4}}-\frac{12M\ell}{b(1-\ell)^{4}}+\frac{12M\ell^{2}}{b(1-\ell)^{4}}-\frac{4M\ell^{3}}{b(1-\ell)^{4}}$
-$\!\!\!\!\!\!+\frac{15\pi M^{2}}{4b^{2}(1-\ell)^{4}}-\frac{3\pi M^{2}\ell}{2b^{2}(1-\ell)^{4}}+\frac{3\pi M^{2}\ell^{3}}{2b^{2}(1-\ell)^{4}}-\frac{3\pi M^{2}\ell^{4}}{4b^{2}(1-\ell)^{4}}$
-$\!\!\!\!\!\!-\frac{3\pi Q^{2}\zeta}{4b^{2}(1-\ell)^{4}e^{\gamma}}+\frac{3\pi Q^{2}\zeta\ell}{4b^{2}(1-\ell)^{4}e^{\gamma}}-\frac{4MQ^{2}\zeta}{3b^{3}(1-\ell)^{4}e^{\gamma}}$
-$\!\!\!\!\!\!+\frac{8MQ^{2}\ell^{2}\zeta}{3b^{3}(1-\ell)^{4}e^{\gamma}}-\frac{4MQ^{2}\ell\zeta}{3b^{3}(1-\ell)^{4}e^{\gamma}}+\mathcal{O}(M^{3},1/b^{4}).$ (30)
-This expression naturally decomposes into three classes of contributions: (i) mass terms proportional to $M/b$ and $M^{2}/b^{2}$, (ii) Kalb-Ramond corrections controlled by powers of $\ell$ and the universal amplification factor $(1-\ell)^{-4}$, and (iii) charge-dependent ModMax terms proportional to $Q^{2}\zeta$, whose magnitude is suppressed by $e^{-\gamma}$ and whose sign is fixed by the branch parameter $\zeta$.
+As noted in Eq.
+(4), the metric function satisfies $f(r)\to 1/(1-\ell)$ as $r\to\infty$.
+As a consequence, the geodesic curvature of the circular boundary $C_{R}$ at large $R$ in the optical geometry evaluates to
+$\lim_{R\to\infty}\kappa(C_{R})\,\frac{dt}{d\phi}=\sqrt{f_{\infty}}=\frac{1}{\sqrt{1-\ell}}\neq 1,$ (26)
+confirming that the optical manifold is not asymptotically Euclidean Jusufi et al.
+(2017); Ono et al.
+(2017).
+Therefore, the standard Gibbons-Werner Gauss-Bonnet formula Gibbons and Werner (2008) must be replaced by the OIA generalisation.
+To make the structure transparent, we define $K\equiv\sqrt{1-\ell}$ and introduce the rescaled metric function
+$h(r)\;\equiv\;K^{2}\,f(r)=1-\frac{2M(1-\ell)}{r}+\frac{\zeta\,Q^{2}\,e^{-\gamma}}{(1-\ell)\,r^{2}},$ (27)
+which by construction satisfies $h(r)\to 1$ as $r\to\infty$.
+Under the coordinate transformation $\tilde{t}=t/K$, $\tilde{r}=K\,r$, the line element (1) takes the Barriola-Vilenkin-type form Jusufi et al.
+(2017)
+$ds^{2}=-\hat{h}(\tilde{r})\,d\tilde{t}^{2}+\frac{d\tilde{r}^{2}}{\hat{h}(\tilde{r})}+\frac{\tilde{r}^{2}}{K^{2}}\,d\Omega^{2},$ (28)
+where
+$\hat{h}(\tilde{r})=1-\frac{2\hat{M}}{\tilde{r}}+\frac{\hat{Q}^{2}}{\tilde{r}^{2}},$ (29)
+with the effective parameters
+$\hat{M}=M(1-\ell)^{3/2},\qquad\hat{Q}^{2}=\zeta\,Q^{2}\,e^{-\gamma}.$ (30)
+The angular part $\tilde{r}^{2}/K^{2}$ in Eq.
+(28) shows that the transformed metric possesses an angular surplus (since $1/K^{2}>1$ for $\ell>0$), in contrast to the angular deficit arising in the Letelier cloud-of-strings or global monopole spacetimes studied in Jusufi et al.
+(2017); Sucu and Sakalli (2025c).
+The physical impact parameter in the rescaled frame is $\tilde{b}=Kb$, where $b$ is the coordinate distance of closest approach in the original frame.
+The rescaled function $\hat{h}$ has exactly the Reissner-Nordström form, so all standard weak-field lensing results can be imported once the OIA boundary correction is included.
+V.2 Gauss-Bonnet method with OIA correction
+The Gauss-Bonnet theorem applied to the non-singular domain $\mathcal{D}_{\infty}$ exterior to the photon trajectory in the optical geometry yields Gibbons and Werner (2008); Jusufi et al.
+(2017); Ono et al.
+(2017)
+$\iint_{\mathcal{D}_{\infty}}\!\mathcal{K}\,dS+\frac{1}{K}\bigl(\pi+\hat{\alpha}\bigr)=\pi,$ (31)
+where $\mathcal{K}$ is the Gaussian curvature and $dS$ is the surface element of the optical manifold, and the factor $1/K$ arises from Eq. (26). Solving for the deflection angle gives
+$\hat{\alpha}=\pi(K-1)-K\iint_{\mathcal{D}_{\infty}}\!\mathcal{K}\,dS.$ (32)
+The first term, $\pi(\sqrt{1-\ell}-1)<0$, is the topological boundary correction absent in the standard Gibbons-Werner approach.
+For the Letelier or global monopole spacetimes, this correction is positive and proportional to the angular deficit Jusufi et al.
+(2017); in the present case the sign is reversed because the KR condensate modifies the lapse rather than the angular metric.
+The Gaussian curvature of the optical metric (25) is determined by Sucu and Sakallı (2025a)
+$\mathcal{K}=-\frac{f^{\prime\,2}-2f\,f^{\prime\prime}}{4},$ (33)
+which in the weak-field region ($r\gg r_{h}$) yields
+$\mathcal{K}\simeq-\frac{2M}{(1-\ell)\,r^{3}}+\frac{3M^{2}+\dfrac{3\zeta\,Q^{2}\,e^{-\gamma}}{(1-\ell)^{3}}}{r^{4}}+\mathcal{O}(r^{-5}).$ (34)
+The leading $r^{-3}$ contribution originates from the mass sector and carries a factor $(1-\ell)^{-1}$ reflecting the renormalized gravitational coupling.
+The $r^{-4}$ terms encode both second-order mass corrections and the charge-dependent ModMax contributions; the latter are suppressed by $e^{-\gamma}$ and amplified by $(1-\ell)^{-3}$.
+To evaluate the curvature integral in Eq.
+(32) to $\mathcal{O}(M^{2}/b^{2})$ accuracy, the straight-line trajectory $r=b/\sin\phi$ is insufficient for the $M^{2}$ coefficient Pereira and others (2025); Araújo Filho and others (2025).
+Instead, one employs the iterative photon trajectory
+$\frac{1}{\tilde{r}}=\frac{\sin\phi}{\tilde{b}}+\frac{\hat{M}\,(1-\cos\phi)^{2}}{\tilde{b}^{2}},$ (35)
+which accounts for the self-consistent bending of the photon path to first post-Minkowskian order Pereira and others (2025).
+With this improved trajectory, the curvature integral yields the correct coefficient $15\pi/(4b^{2})$ for the $M^{2}$ contribution, replacing the lower value $3\pi/(4b^{2})$ obtained from the straight-line approximation.
+V.3 Weak-field deflection angle and branch-dependent signatures
+Carrying out the integration in Eq. (32) with the iterative trajectory (35) and the Gaussian curvature (34), we obtain the vacuum deflection angle for the Kalb-Ramond-ModMax black hole:
+$\hat{\alpha}_{\rm vac}$ $=$ $\pi\!\left(\sqrt{1-\ell}\,-1\right)+\frac{4M(1-\ell)}{b}$ (36)
+$\!\!\!\!\!\!+\frac{\pi\!\left[15M^{2}(1-\ell)^{2}-\dfrac{3\zeta\,Q^{2}\,e^{-\gamma}}{1-\ell}\right]}{4b^{2}}$
+$\!\!\!\!\!\!-\frac{4M\zeta\,Q^{2}\,e^{-\gamma}}{3b^{3}}+\mathcal{O}(b^{-4}).$
+This expression naturally decomposes into four physically distinct contributions: (i) the topological boundary term $\pi(\sqrt{1-\ell}-1)<0$, which arises from the non-Euclidean asymptotic structure and reduces the total deflection; (ii) the leading mass term $4M(1-\ell)/b$, which is suppressed relative to the Schwarzschild value $4M/b$; (iii) second-order terms at $\mathcal{O}(b^{-2})$ encoding gravitational self-interaction and charge-dependent ModMax corrections; and (iv) a mixed mass-charge term at $\mathcal{O}(b^{-3})$, which is independent of $\ell$ and coincides with the standard Reissner-Nordström expression.
+Expanding to first order in $\ell$, which is the physically relevant regime given observational bounds on Lorentz-violating parameters Masood (2024), the deflection angle simplifies to
+$\hat{\alpha}_{\rm vac}$ $\approx$ $-\frac{\pi\,\ell}{2}+\frac{4M}{b}\left(1-\ell\right)$ (37)
+$+\frac{\pi\!\left(15M^{2}-3\zeta\,Q^{2}\,e^{-\gamma}\right)}{4b^{2}}+\mathcal{O}(\ell^{2},\ell/b^{2}),$
+where the $\ell$-dependence enters exclusively through the boundary term and the leading mass contribution.
 Several physically relevant limits follow immediately. In the Schwarzschild limit $(\ell\to 0,\gamma\to 0,Q\to 0)$, Eq. (36) reduces to
-$\beta_{\rm Sch}=\frac{4M}{b}+\frac{15\pi M^{2}}{4b^{2}}+\mathcal{O}(b^{-3}),$ (31)
-recovering the standard general-relativistic result Epstein and Shapiro (1980); Weinberg (2013).
+$\hat{\alpha}_{\rm Sch}=\frac{4M}{b}+\frac{15\pi M^{2}}{4b^{2}}+\mathcal{O}(b^{-3}),$ (38)
+recovering the standard general-relativistic result including the correct second-order coefficient Epstein and Shapiro (1980); Weinberg (2013); Pereira and others (2025).
 For the ordinary branch $(\zeta=+1)$ with $\ell\to 0$ and $\gamma\to 0$, the expression reduces to the Reissner-Nordström deflection angle, i.e.
-$\beta_{\rm RN}=\frac{4M}{b}+\frac{15\pi M^{2}}{4b^{2}}-\frac{3\pi Q^{2}}{4b^{2}}-\frac{4MQ^{2}}{3b^{3}}+\mathcal{O}(b^{-4}),$ (32)
+$\hat{\alpha}_{\rm RN}=\frac{4M}{b}+\frac{15\pi M^{2}}{4b^{2}}-\frac{3\pi Q^{2}}{4b^{2}}-\frac{4MQ^{2}}{3b^{3}}+\mathcal{O}(b^{-4}),$ (39)
 where the electromagnetic contribution partially counteracts gravitational focusing.
-On the other hand, in the phantom branch $(\zeta=-1)$ the sign reversal of the charge sector transforms the electromagnetic contribution into an additional focusing mechanism.
-Even in the limit $\ell\to 0$ and $\gamma\to 0$, the deflection angle becomes
-$\beta_{\rm phantom}=\frac{4M}{b}+\frac{15\pi M^{2}}{4b^{2}}+\frac{3\pi Q^{2}}{4b^{2}}+\frac{4MQ^{2}}{3b^{3}}+\mathcal{O}(b^{-4}),$ (33)
-which has no analogue in standard general relativity. In this sector, both mass and charge contribute constructively to light bending, leading to systematically larger deflection angles.
+In the phantom branch $(\zeta=-1)$ with $\ell\to 0$ and $\gamma\to 0$, the sign reversal of the charge sector transforms the electromagnetic contribution into an additional focusing mechanism:
+$\hat{\alpha}_{\rm phantom}=\frac{4M}{b}+\frac{15\pi M^{2}}{4b^{2}}+\frac{3\pi Q^{2}}{4b^{2}}+\frac{4MQ^{2}}{3b^{3}}+\mathcal{O}(b^{-4}),$ (40)
+where both mass and charge contribute constructively to light bending, producing systematically larger deflection angles.
 (a) Ordinary branch ($\zeta=+1$).
 (b) Phantom branch ($\zeta=-1$).
-Figure 8: Vacuum deflection angle $\beta_{\rm vac}$ as a function of the impact parameter $b$ for KR-ModMax black holes with fixed $M=1$, $Q=0.5$, and $\gamma=2$.
-The color scale indicates the Kalb-Ramond parameter $\ell\in[0,0.5]$.
-Panel (a) shows the ordinary branch ($\zeta=+1$), where the electromagnetic contribution partially counteracts gravitational focusing, leading to suppressed deflection at small $b$.
-Increasing $\ell$ enhances the mass-dominated terms through the $(1-\ell)^{-4}$ amplification.
-Panel (b) displays the phantom branch ($\zeta=-1$), where the sign reversal of the electromagnetic sector transforms the charge contribution into an additional focusing mechanism.
-Consequently, the deflection angle remains strictly positive and is notably enhanced at small $b$, exceeding the ordinary-branch values by up to an order of magnitude as $\ell$ increases.
-This qualitative distinction serves as an observational discriminator between the two electromagnetic sectors.
-These qualitative differences are evident in the numerical results shown in Fig.
-8, where the deflection angle is plotted for varying Kalb-Ramond parameter $\ell\in[0,0.5]$ at fixed $\gamma=2$.
-In the ordinary branch, the competition between mass-induced focusing and charge-induced defocusing produces smaller deflection angles, particularly at low $\ell$.
-As $\ell$ increases, the universal amplification factor $(1-\ell)^{-4}$ enhances all contributions, leading to progressively larger deflection at all impact parameters.
-In the phantom branch, the deflection angle remains strictly positive and is substantially enhanced relative to the ordinary branch, exceeding the latter by more than an order of magnitude at small $b$ for larger $\ell$ values.
-This pronounced enhancement provides a clear geometric manifestation of the branch dichotomy.
-From an observational perspective, these branch-dependent signatures translate into potentially measurable effects.
-Phantom configurations would produce significantly larger Einstein ring radii and enhanced magnification in strong-lensing systems, while the $\ell$-dependence of the weak-field deflection could imprint small but systematic astrometric shifts in high-precision measurements of stellar orbits near supermassive black holes.
-Together with the thermodynamic and photon-sphere properties discussed in previous sections, gravitational lensing offers a complementary channel for discriminating between the ordinary and phantom Kalb-Ramond-ModMax black hole branches.
+Figure 8: Vacuum deflection angle $\hat{\alpha}_{\rm vac}$ as a function of the impact parameter $b$ for KR-ModMax black holes with fixed $M=1$, $Q=0.5$, and $\gamma=2$.
+The colour scale indicates the Kalb-Ramond parameter $\ell\in[0,0.5]$.
+Panel (a) shows the ordinary branch ($\zeta=+1$), where the electromagnetic contribution partially counteracts gravitational focusing.
+Increasing $\ell$ reduces the total deflection through the negative boundary term $\pi(\sqrt{1-\ell}-1)$ and the suppressed effective mass $M_{\rm eff}=M(1-\ell)$.
+Panel (b) displays the phantom branch ($\zeta=-1$), where the sign reversal of the electromagnetic sector enhances focusing.
+Although the phantom-branch deflection angles remain larger than their ordinary counterparts, increasing $\ell$ reduces both branches relative to their Schwarzschild baselines through the universal topological correction.
+The qualitative features of Eq.
+(36) are illustrated in Fig.
+8.
+A distinctive prediction of the OIA-corrected analysis is that the Kalb-Ramond parameter $\ell$ reduces the deflection angle through two mechanisms: the negative topological boundary contribution $\pi(\sqrt{1-\ell}-1)$, and the suppression of the effective gravitational mass to $M(1-\ell)$ in the leading term.
+This contrasts qualitatively with the Letelier cloud-of-strings spacetime, where the analogous topological term is positive and enhances deflection Jusufi et al.
+(2017).
+The difference is rooted in the distinct locations of non-flatness: the Letelier string cloud modifies the angular metric ($g_{\phi\phi}=(1-A)r^{2}$), producing an angular deficit, while the KR condensate modifies the lapse ($f_{\infty}=1/(1-\ell)$), producing an effective angular surplus in the optical geometry.
+The branch dichotomy persists: the phantom branch ($\zeta=-1$) yields larger deflection angles than the ordinary branch ($\zeta=+1$) at fixed $\ell$, due to the constructive alignment of mass and charge contributions.
+However, both branches experience the same $\ell$-dependent reduction, so the ratio of phantom-to-ordinary deflection angles remains sensitive to the charge-to-mass ratio and the ModMax parameter $\gamma$.
+From an observational perspective, the reduction of the deflection angle with increasing $\ell$ would manifest as smaller Einstein ring radii and decreased magnification in strong-lensing systems compared to the Schwarzschild prediction at the same mass, providing a distinctive observational signature of Kalb-Ramond Lorentz symmetry breaking.
+Together with the thermodynamic and photon-sphere properties discussed in previous sections, gravitational lensing offers a complementary channel for constraining the KR parameter $\ell$ and discriminating between the ordinary and phantom Kalb-Ramond-ModMax black hole branches.
 VI Plasma effects on photon propagation
@@
 Optical properties provide an independent and complementary window into the branch structure.
-Using the Gauss-Bonnet approach to weak gravitational lensing, we found that the ordinary branch generally produces reduced deflection angles compared to the Schwarzschild case, reflecting the partially repulsive role of the electromagnetic sector.
-In contrast, the phantom branch yields systematically enhanced deflection, since the sign-reversed electromagnetic contribution reinforces gravitational focusing.
+Since the KR-ModMax metric satisfies $f(r)\to 1/(1-\ell)\neq 1$ at spatial infinity, the standard Gibbons-Werner formulation of the Gauss-Bonnet theorem is inapplicable.
+We therefore employed the Ono-Ishihara-Asada extension, which introduces a topological boundary correction $\pi(\sqrt{1-\ell}-1)<0$ that universally reduces light bending relative to the Schwarzschild baseline.
+The effective gravitational mass $M(1-\ell)$ in the leading term further suppresses deflection, a qualitative departure from the Letelier cloud-of-strings spacetime, where the analogous topological term is positive and enhances light bending.
+This sign reversal reflects the distinct origin of non-flatness: the KR condensate modifies the lapse rather than the angular metric, producing an effective angular surplus in the optical geometry.
+The branch dichotomy $\zeta=\pm 1$ persists through the charge-dependent second-order terms, with the phantom branch yielding larger deflection angles due to the constructive alignment of mass and charge contributions.
 This distinction persists, in a modified form, when photon propagation is studied in plasma environments.